### v1

# Efficient Retrieval with Learned Similarities

Abstract
Retrieval plays a fundamental role in recommendation systems, search, and natural language processing by efficiently finding relevant items from a large corpus given a query. Dot products have been widely used as the similarity function in such retrieval tasks, thanks to Maximum Inner Product Search (MIPS) that enabled efficient retrieval based on dot products. However, state-of-the-art retrieval algorithms have migrated to learned similarities. Such algorithms vary in form; the queries can be represented with multiple embeddings, complex neural networks can be deployed, the item ids can be decoded directly from queries using beam search, and multiple approaches can be combined in hybrid solutions. Unfortunately, we lack efficient solutions for retrieval in these state-of-the-art setups. Our work investigates techniques for approximate nearest neighbor search with learned similarity functions. We first prove that Mixture-of-Logits (MoL) is a universal approximator, and can express all learned similarity functions. We next propose techniques to retrieve the approximate top $K$ results using MoL with a tight bound. We finally compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
1 Introduction
Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors. Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022). In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions. Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020). As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf. More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups (Tay et al., 2022; Wang et al., 2022). Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021). Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
Supporting efficient retrieval with these diverse learned similarities is challenging. Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018; Zhai et al., 2023), while large scale recommendation systems commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019) with low end-to-end latency. When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023). What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface for these techniques makes it even more difficult for designing a general solution for efficient retrieval.
Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage. Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them. Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$). This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022), and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Gao et al., 2021; Zhai et al., 2024).
This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL. To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities. We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions. We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound. Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014). Importantly, our approach efficiently utilizes modern accelerators like GPUs due to higher arithmetic intensity. We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
2 Expressiveness of Mixture of Logits
In this section, we describe Mixture of Logits (MoL) and show that MoL is expressive enough to represent any learned similarity function. Table 1 summarizes the notations in this paper.
We first describe the learned similarity function with Mixture of Logits (MoL).
Notation Description
$q$ ($Q$, $|Q|$) query (set of queries, number of queries)
$x$ ($X$, $|X|$) item (set of items, number of items)
$\phi(q,x)$ the learned similarity function, i.e., Mixture-of-Logits (MoL).
$P$ ($P_{q}$, $P_{x}$) number of component-level embeddings. With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of component-level embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
$\pi_{p}(q,x)$ ($\pi_{p_{q},p_{x}}(q,x)$) weight for the $p$-th (or $p_{q}$-th by $p_{x}$-th with outer product) embedding set for $(q,x)$.
$f(q)$ ($f_{p}(q)$) learned embedding for the query ($p$-th component-level query embedding)
$g(x)$ ($g_{p}(x)$) learned embedding for the item ($p$-th component-level item embedding)
$\langle f(q),g(x)\rangle$ the dot product similarity function; $\langle f(q),g(x)\rangle=g(x)^{T}f(q)$.
Table 1: Table of Notations.
Mixture of Logits (MoL)
MoL applies an adaptive gating weight $\pi_{p}(q,x)$ to $P$ groups of component-level dot products $\langle f_{p}(q),g_{p}(x)\rangle$. Following Zhai et al. (2023):
$\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ (1)
To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings (where $P_{q}\times P_{x}=P$) and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
$\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$ (2)
We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper as the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
Figure 1: Mixture-of-logits (MoL) learned similarity.
Now we show that any high rank matrix can be decomposed into a mixture of logits with low rank matrices. Specifically, without loss of generality, we prove the following:
Theorem 1.
[…]
First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$:
Lemma 1.
MoL decomposition when $Rank(A)\leq d$: Let $A$ be a matrix as defined in Theorem 1. If the rank of $A$ is less than or equal to $d$, then we have $A=\pi\circ A$, where $\pi(i,j)=1$ for any $1\leq i\leq n,1\leq j\leq m$.
Then we prove for the case where the rank of $A$ is greater than $d$. Without loss of generality, we prove the case where the matrix has full rank, i.e., $Rank(A)=n$:
Lemma 2.
MoL decomposition when $Rank(A)=n$: Let $A$ be a matrix as defined in Theorem 1. Then there exists $\pi,B_{1},B_{2}$ such that $|A-(\pi\circ B_{1}+(1-\pi)\circ B_{2})|<\epsilon$, where $Rank(B_{1})\leq d$, $Rank(B_{2})\leq d$, and $0\leq\pi(i,j)\leq 1$ for $1\leq i\leq n,1\leq j\leq m$.
Proof.
Because $A$ is a matrix of rank $n$, it can be rewritten as $A=UI_{n}V$, where $I_{n}$ is an identity matrix with rank $n$. Thus, $A_{ij}=\sum_{k=1}^{n}U_{ik}V_{kj},1\leq i\leq n,1\leq j\leq m$. Let $A^{\prime}$ be a matrix of $n\times m$, where $A^{\prime}_{ij}=\lambda_{ij}\cdot\sum_{k=1}^{d}U_{ik}V_{kj}$ for $1\leq i\leq n,1\leq j\leq m$. Here, $\lambda_{ij}=1+\frac{\sum_{k=d+1}^{n}U_{ik}V_{kj}}{\sum_{k=1}^{d}U_{ik}V_{kj}}$ if $\sum_{k=1}^{d}U_{ik}V_{kj}\neq 0$, otherwise $\lambda_{ij}=1+\frac{\sum_{k=d+1}^{n}U_{ik}V_{kj}}{\epsilon}$. Thus, we have $|A-A^{\prime}|\leq\epsilon$.
Let $\lambda_{min}=\min{\lambda_{ij}}$, and $\lambda_{max}=\max{\lambda_{ij}}$. Let $B_{1}=\lambda_{min}UI_{d}V$, $B_{2}=\lambda_{max}UI_{d}V$, where $I_{d}$ is the identity matrix of rank $d$. We have $A^{\prime}_{ij}=\lambda_{ij}\sum_{k=1}^{d}U_{ik}V_{kj}=\pi(i,j)\cdot B_{1ij}+(1-\pi(i,j))\cdot B_{2ij}$, where $\pi(i,j)=\frac{\lambda_{max}-\lambda_{ij}}{\lambda_{max}-\lambda_{min}}$. Because $\lambda_{min}\leq\lambda_{ij}\leq\lambda_{max}$, we have $0\leq\pi(i,j)\leq 1$.
Thus, we have constructed $B_{1},B_{2},\pi$ such that $|A-(\pi\circ B_{1}+(1-\pi)\circ B_{2})|=|A-A^{\prime}|\leq\epsilon$. ∎
Remark Here, we have shown that any high-rank matrix can be expressed as a mixture of logits of two low-rank matrices. Note that our decomposition is not intended to be used as a distillation of the original high-rank matrix. For instance, it is prohibitively expensive to populate the full matrix with a learned similarity function. In addition, the proof also does not indicate that having a sum of two mixture of logits is sufficient to train the embeddings and the learned similarity function. It is well-known that overparameterization is often necessary to enable efficient and performant training.
3 Retrieval Algorithms
In this section, we describe the problem of retrieving the top $K$ items with MoL and both exact and approximate algorithms for retrieval. Formally, we define the top $K$ retrieval problem as follows:
Definition 1.
Top $K$ with MoL: Let $q$ be a query and $X$ be a set of items, where both the query $q$ and each item $x\in X$ are associated with $P$ embeddings. Together we have $P$ pairs of embeddings, $(f_{p}(q),g_{p}(x)),1\leq p\leq P$. Let $\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ be the similarity score of $q,x$, where $x\in X$. The top $K$ query with MoL returns the $K$ items from $X$ with the highest $\phi(q,x)$s.
For approximate top $K$ retrieval with MoL, we define the gap of the approximate and exact top $K$ results as follows:
Definition 2.
Gap of approximate top $K$: Let $q$ be a query and $X_{K}$ be the set of exact top $K$ items for the query $q$ from a set of items $X$. Let $X^{*}$ be the approximate top $K$ results, where $X^{*}\subseteq X$. Let $S=\min\{\phi(q,x),x\in X^{*}\}$ and $S^{\prime}=\max\{\phi(q,x),x\in X_{K}\setminus X^{*}\}$. We call $S_{\Delta}=S^{\prime}-S$ the gap of the top $K$ with $X^{*}$.
3.1 Exact algorithm
The brute-force algorithm to retrieve the exact top $K$ with MoL is to evaluate $\phi(q,x)$ for each query $q$ and item $x$. This algorithm can be prohibitively expensive if the number of items is large. Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1.
Algorithm 1 Exact top $K$ algorithm.
1:query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
2:exact top $K$ items
3:$G\leftarrow\emptyset$
4:for $p\in P$ do
5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings
7:$S_{min}\leftarrow\infty$
8:for $x\in G$ do
9:    $s\leftarrow MoL(q,x)$
10:    if $s<S_{min}$ then $S_{min}\leftarrow s$
11:$G^{\prime}\leftarrow\emptyset$
12:for $p\in P$ do
13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items with a dot product higher than or equal to $S_{min}$ for $f_{p}(q)$ from $X_{p}$
14:return $BruteForceTopKMoL(q,G^{\prime})$$\triangleright$ Retrieve the top $K$ items from $G^{\prime}$ with MoL
We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4). Then we evaluate the MoL score of the items in the set $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8). Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11). Finally, we evaluate the MoL scores of the items in $G^{\prime}$ exhaustively, and return the top $K$ items with the highest scores (line 12).
We argue that Algorithm 1 retrieves the exact top $K$ items with MoL. Let $X_{K}$ be the set of the exact top $K$ items and $X^{\prime}$ be the result of Algorithm 1. Let $x\in X_{K}$ and $\phi(q,x)$ be the MoL score of $x$ and $q$. Since $x$ has the highest top $K$ score with MoL, $\phi(q,x)\geq S_{min}$. Since the MoL score is a weighted score over the dot product scores, we have $\max\{\langle f_{p}(q),g_{p}(x)\rangle,1\leq p\leq P\}\geq\phi(q,x)\geq S_{min}$. Since Algorithm 1 retrieves all the items with a dot product higher than or equal to $S_{min}$ of $q$ for each embedding $q_{p}$ (line 9-11), we have $x\in G^{\prime}$. Thus, $x\in X^{\prime}$. So we have shown that $X_{K}=X^{\prime}$.
3.2 Approximate algorithms
In the exact algorithm shown in Algorithm 1, we need to retrieve all the items with a dot product higher than or equal to a threshold. When the threshold is a loose filter of the item set, which may happen when the dot product scores are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large set of candidate items can still be expensive. Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact top $K$ algorithm.
In both heuristics, we perform a two-stage retrieval as shown in Algorithm 2. In the first stage, we retrieve a set of $K^{\prime}$ candidate items that are potentially high in MoL score by using dot products (line 2). Note that $K^{\prime}$ can be larger than $K$, e.g., due to oversampling. In the second stage, we evaluate the MoL score of the candidate items and return the top $K$ items (line 3).
Algorithm 2 Approximate top $K$ algorithms.
1:a query $q$, a set of items $X$
2:approximate top $K$ items
3:function ApproxTopK($q,X,K,K^{\prime}$)
4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidate items
5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items with MoL
6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
7:union of top $K$ items over $P$ component-level embeddings by dot product
8:function TopKPerEmbedding($q,X,K$)
9:    $G\leftarrow\emptyset$
10:    for $p\in P$ do
11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product
13:    return $G$
14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embedding $f_{p}(q),g_{p}(x)$
15:top $K$ items based on the averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
16:function TopKAvg($q,X,K$)
17:    $q^{\prime}\leftarrow\sum_{p=1}^{P}f_{p}(q)$
18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed and materialized.
19:    return $TopKDotProduct(q^{\prime},X^{\prime},K)$
Here, we describe two heuristics to retrieve the candidate item set:
Top $K$ per embedding.
Given a query $q$ and a set of items $X$, for each embedding set $p$, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$). Return the union across $P$ queries.
The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding by dot product. We analyze the gap of this heuristic as follows:
Theorem 2.
Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\phi(q,x),x\in X_{K+1,p}\}$. Let $S_{min}$ be the top $K^{th}$ MoL score of the items in $\cup_{p}X_{K,p}$, then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Note that there exists an MoL such that $S_{\Delta}=S-S_{min}$, i.e., when $\pi_{p}(q,x)=1$ for $x_{p}=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p}\}$. Thus, the upper bound of $S_{\Delta}$ is tight.
Top $K$ average.
Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
Note that the top $K$ average heuristic returns the exact top $K$ items when the gating weight distribution in MoL, $\pi$, is uniform.
This heuristic is interesting for two reasons. First, the items retrieved by this heuristic are likely to be the top $K$ items of MoL when the weight distribution is more balanced. This complements the heuristic that retrieves top $K$ per embedding. Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding sets $P=P_{q}\times P_{x}$.
Formally, let $P=P_{q}\cdot P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$. The average dot product can be computed as
$\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}=\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}=\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$ (3)
Thus, we can preprocess the embeddings of the items and the query, so the number of embeddings accessed is $1$ per item for a given query, regardless of the overall number of component-level embeddings used by MoL, i.e., $P$.
Finally, we can combine the candidates retrieved from top $K$ per embedding group and the top $K$ average as the following:
Combined top $K$.
Given a query $q$, a set of items $X$, and $K$, return the union of the items from the top $K$ per embedding group across the $P$ groups and the top $K$ items from the top $K$ average.
Theorem 3.
Upper bound of the gap of combined top $K$. Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{min}$ as defined in Theorem 2. Let $X^{\prime}_{K}$ be the top $K$ items from top $K$ average. Let $S^{\prime}=\max\{\phi(q,x),x\in X\setminus(\cup_{p}X_{K,p}\cup X^{\prime}_{K})\}$. Then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
Remark Similar to Theorem 2, the upper bound of the gap is tight. In practice, we can configure the $K$ to be different for the two heuristics, i.e., $K_{1}$ and $K_{2}$. For example, when the weight distribution $\pi$ is more balanced, $K_{2}$ can be configured to be larger as the top $K$ average approach approximates MoL well while being more computationally efficient.
4 Evaluation
In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3. Our code and pretrained model checkpoints are publicly available at https://github.com/bailuding/rails.
4.1 Setup
Workloads.
We consider a classical retrieval task (next-item prediction) in recommendation systems (Covington et al., 2016; Zhu et al., 2018; Li et al., 2019; Zhai et al., 2023). We employ three widely used datasets, MovieLens (Harper and Konstan, 2015) (the 1M and 20M subsets), and Amazon Reviews (McAuley et al., 2015) (the largest Books subset). We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets (Kang and McAuley, 2018; Zhai et al., 2023). In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step. Table 2 shows the statistics of these three workloads.
Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ component-level embedding dimension
ML-1M 6,040 3,649 8 4 64
ML-20M 138,493 24,186 8 4 128
Books 694,897 674,044 8 8 32
Table 2: Workload statistics.
Model training.
We implemented MoL on top of two popular self-attention sequential encoder architectures, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024). The models are implemented in PyTorch and are trained with 2x NVIDIA RTX 6000 Ada GPUs.
Method HR@1 HR@10 HR@50 HR@200 MRR
ML-1M SASRec .0610 .2806 .5437 .7498 .1307
SASRec + MoL .0672 .3019 .5711 .7650 .1431
HSTU .0774 .3306 .5897 .7832 .1571
HSTU + MoL .0875 .3412 .6013 .7877 .1692
ML-20M SASRec .0674 .2889 .5503 .7661 .1388
SASRec + MoL .0731 .3068 .5645 .7768 .1498
HSTU .0965 .3556 .6143 .8074 .1800
HSTU + MoL .1014 .3661 .6234 .8116 .1876
Books SASRec .0058 .0306 .0754 .1431 .0153
SASRec + MoL .0083 .0402 .0970 .1813 .0203
HSTU .0101 .0469 .1066 .1876 .0233
HSTU + MoL .0150 .0613 .1292 .2167 .0315
Table 3: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
Metrics.
We use Recall (Hit Rate) as the main metric. It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3). Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 21.4% in HR@1, an average of 13.7% in HR@10, and an average of 16.2% in MRR across the six scenarios, validating the importance of learned similarity functions. We evaluate our proposed retrieval algorithms relative to the best performing rows, specifically “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the following section.
4.2 Top $K$ retrieval performance
Method HR@1 HR@5 HR@10 HR@50 HR@100 Latency / ms
ML-1M BruteForce 1.00 1.00 1.00 1.00 1.00 .64$\pm$.03
TopKPerEmbd5 .838 .768 .681 .499 .437 .95$\pm$.03
TopKPerEmbd10 .932 .905 .842 .678 .618 .94$\pm$.05
TopKPerEmbd50 1.00 .997 .999 .977 .954 .95$\pm$.05
TopKPerEmbd100 1.00 1.00 1.00 .995 .990 1.24$\pm$.04
TopKAvg200 .980 .970 .967 .958 .948 .74$\pm$.06
TopKAvg500 .992 .986 .989 .990 .990 .71$\pm$.04
TopKAvg1000 .996 .994 .998 .997 .999 .72$\pm$.03
CombTopK5_200 1.00 .997 .988 .975 .960 1.09$\pm$.08
CombTopK50_500 1.00 .999 .999 .998 .997 1.13$\pm$.04
ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 3.08$\pm$.02
TopKPerEmbd5 .755 .702 .645 .467 .402 1.19$\pm$.05
TopKPerEmbd10 .946 .876 .817 .645 .564 1.17$\pm$.05
TopKPerEmbd50 .993 .986 .977 .931 .900 1.55$\pm$.03
TopKPerEmbd100 .998 .997 .994 .979 .966 2.26$\pm$.03
TopKAvg200 .992 .992 .992 .980 .958 .76$\pm$.04
TopKAvg500 .992 .993 .995 .995 .993 .77$\pm$.04
TopKAvg1000 .992 .994 .996 .998 .997 .75$\pm$.03
CombTopK5_200 1.00 .998 .998 .983 .962 1.32$\pm$.06
CombTopK50_500 1.00 1.00 1.00 .999 .996 1.85$\pm$.03
Books BruteForce 1.00 1.00 1.00 1.00 1.00 149.02$\pm$.49
TopKPerEmbd5 .907 .915 .809 .509 .396 20.17$\pm$.06
TopKPerEmbd10 .936 .986 .914 .698 .561 20.33$\pm$.06
TopKPerEmbd50 .993 .992 .994 .956 .902 21.09$\pm$.04
TopKPerEmbd100 1.00 .995 .996 .985 .959 22.53$\pm$.06
TopKAvg200 1.00 .978 .948 .845 .767 .68$\pm$.03
TopKAvg500 1.00 .992 .996 .919 .875 .68$\pm$.03
TopKAvg1000 1.00 1.00 .996 .963 .939 .84$\pm$.01
TopKAvg2000 1.00 1.00 1.00 .980 .968 1.04$\pm$.02
TopKAvg4000 1.00 1.00 1.00 .996 .987 1.64$\pm$.04
CombTopK5_200 .979 .984 .981 .892 .818 20.29$\pm$.05
CombTopK50_500 .993 .989 .994 .975 .961 21.57$\pm$.06
CombTopK100_1000 1.00 .997 .996 .993 .992 23.00$\pm$.08
Table 4: Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency with standard deviation measured over the evaluation of retrieval for a batch of queries (where the batch size is 32). (Relative) hit rate higher than .99 is marked in bold.
We evaluate the following methods on their top $K$ retrieval performance:
  • Brute-force top $K$ (BruteForce): Evaluate the MoL scores for all items and return the top $K$ items. This is the ground truth in our top $K$ evaluation  [^1] We omit the baseline with the two-pass exact algorithm (Section 3.1) because the range-based item retrieval can still be expensive when the range threshold is loose. Empirically, the brute-force top $K$ is more efficient on our datasets. We leave the efficient implementation of the two-pass exact algorithm as future work..
  • Per embedding top $K$ (TopKPerEmbd($N$)): This algorithm is described in Section 3.2. $N$ is the number of candidate items retrieved from each embedding set, where $N\times P\geq K$.
  • Average top $K$ (TopKAvg($N$)): This algorithm is described in Section 3.2. $N$ is the number of the candidate items retrieved by average dot products, where $N\geq K$.
[…]
Table 4 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023). We set batch size to $32$ for ML-1M, ML-20M, and Books. We perform evaluation on a single NVIDIA RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.
We observe that our approximate heuristics achieve high HR with oversampling. For example, TopKAvg500 is >.$>.99$ in relative HR across the board for ML-1M and ML-20M, and TopKAvg4000 is >.$>.99$ in relative HR across the board for Books. In addition, the combined top $K$ algorithm can outperform both TopKPerEmbd and TopKAvg of the corresponding configurations, sometimes significantly, e.g., CombTopK5_200 vs. TopKPerEmbd5 and TopKAvg200 on Books. This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions vary in MoL.
In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baseline, especially as the number of items in the dataset becomes large. For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of 4×4$\times$4 × and 91×91$\times$91 × in latency for ML-20M and Books, respectively. While the latency of the retrieval grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods. For example, the latency of retrieval increases by 4.8×4.8$\times$4.8 × from ML-1M to ML-20M in BruteForce, while the growth rate is 1.8×1.8$\times$1.8 × and $\times1$ for TopKPerEmbd100 and TopKAvg1000 respectively. Thus, we expect that the speedup of the approximate methods to become even more prominent as the size of the dataset further increases.
We also notice that TopKAvg tends to be more efficient than TopKPerEmbd with comparable HR, e.g., TopKAvg2000 vs. TopKPerEmbd100 on Books with $\times2$ speedup in latency. We believe that this is mainly due to two reasons. First, when the HR is comparable, the maximal number of candidate items from TopKPerEmbd is larger than that of TopKAvg. Second, compared to TopKPerEmbd, the computation of TopKAvg is agnostic to the number of embedding sets because of the materialization optimization described in Section 3.2. Interestingly, we also see that the combined top $K$ is more efficient than the summation of the latency of its individual components, e.g., CombTopK5_200 is 1.5×1.5$\times$1.5 × faster than the sum of the latency from TopKPerEmbd5 and TopKAvg200 on ML-20M. This is because our implementation reduces the overhead of the combined method by consolidating some shared processing of the two components.
Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ can be used if the target HR is extremely stringent.
Limitations
[…]
5 Related work
Retrieval in Recommendation Models.
Most modern recommendation models follow a classical two-stage information retrieval paradigm (Covington et al., 2016), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019) are first filtered down to hundreds to thousands in the retrieval or candidate generation stage, followed by the ranking stage that produces the final top 1-10 candidates. Earlier work on scaling large-scale recommendation models primarily utilize dual-encoder setups(Covington et al., 2016) or sometimes graph algorithms (Eksombatchai et al., 2018). Researchers quickly realized that dot products limited retrieval stage’s performance, and introduced multi-query embeddings (Li et al., 2019) and specialized neural networks (He et al., 2017) to alleviate the expressiveness constraints in retrieval. Learned index structures, or generative retrieval, map each candidate item to a series of tokens and leverage beam search to replace approximate nearest neighbor search. These tokens are represented either implicitly during tree traversal (Zhu et al., 2018; Zhuo et al., 2020) or explicitly (Gao et al., 2021). However, they introduce stability challenges as recommendation systems need to support billion-scale dynamic realtime updated set of items. A latest trend is to exploit characteristics of GPUs, TPUs, and other accelerators; learned similarity functions as well as hardware-aware algorithms have been shown to achieve significant quality improvements (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
Efficient Nearest Neighbor Search (NNS).
Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees while maintaining max-heap property have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020). Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize GPUs; to address this issue, accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024).
6 Conclusion
We have analyzed techniques for efficient retrieval with learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and empirically MoL consistently outperforms dot products in HR across 6 scenarios by an average of 21.4% in HR@1 and 13.7% in HR@10, setting new state-of-the-art on recommendation datasets. We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds. Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. We believe this result is of interest to both recommendation system research communities and researchers working on enabling advanced learned similarities in vector databases.

### v2

@@
 Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors.
-Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Lewis et al., 2020; Borgeaud et al., 2022).
-In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
-Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions.
-Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020).
-As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf.
-More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups (Tay et al., 2022; Wang et al., 2022).
-Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021).
-Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
+Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items [5, 9, 26, 25, 2].
+In many concrete use cases, such as vector databases [19], the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
+Despite the popularity of dot products and numerous work done to improve their efficiency [18, 35, 28, 4], state-of-the-art retrieval algorithms have long moved to various learned similarity functions.
+Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values [26, 22].
+As another example, Probabilistic Label Trees (PLTs) [16] and Tree-based Deep Models (TDMs) [45, 46] map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf.
+More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups [38, 40].
+Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products [8].
+Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms [14, 41, 34, 42].
 Supporting efficient retrieval with these diverse learned similarities is challenging.
-Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018; Zhai et al., 2023), while large scale recommendation systems commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019) with low end-to-end latency.
-When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023).
-What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface for these techniques makes it even more difficult for designing a general solution for efficient retrieval.
+Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps [45, 42], while recommendation and NLP deployments commonly need to handle billions of items [6, 26, 3] with low end-to-end latency.
+When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force [14] or heuristics [42].
+What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult for designing a general solution for efficient retrieval.
 Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage.
@@
 Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$).
-This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022), and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Gao et al., 2021; Zhai et al., 2024).
+This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases [19, 4], and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data [5, 26, 8, 43].
 This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL.
@@
 We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions.
+Compared with prior work on learned similarities [14, 38] which often fail to outperform dot product baselines in large-scale settings [33, 36], our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on corpus with hundreds of millions to billions of items [42, 3].
 We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound.
-Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014).
-Importantly, our approach efficiently utilizes modern accelerators like GPUs due to higher arithmetic intensity.
+Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS [18, 35, 19, 12].
 We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
+Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity [42], which results in MIPS-level inference latency and throughput.
+This concludes both theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
 2 Expressiveness of Mixture of Logits
@@
 $\phi(q,x)$ the learned similarity function, i.e., Mixture-of-Logits (MoL).
-$P$ ($P_{q}$, $P_{x}$) number of component-level embeddings.
-With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of component-level embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
+$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ low-rank embeddings ("component-level embeddings") to represent $q$ and $x$.
+With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
 $\pi_{p}(q,x)$ ($\pi_{p_{q},p_{x}}(q,x)$) weight for the $p$-th (or $p_{q}$-th by $p_{x}$-th with outer product) embedding set for $(q,x)$.
@@
 $g(x)$ ($g_{p}(x)$) learned embedding for the item ($p$-th component-level item embedding)
+$d_{P}$ dimensionality of low-rank (component-level) embeddings. $f_{p}(q),g_{p}(q)\in\mathbb{R}^{d_{P}}$.
 $\langle f(q),g(x)\rangle$ the dot product similarity function; $\langle f(q),g(x)\rangle=g(x)^{T}f(q)$.
@@
 Mixture of Logits (MoL)
-MoL applies an adaptive gating weight $\pi_{p}(q,x)$ to $P$ groups of component-level dot products $\langle f_{p}(q),g_{p}(x)\rangle$. Following Zhai et al. (2023):
+MoL [42, 3] assumes that the query $q$ and the item $x$ are already mapped to $P$ groups of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively.
+MoL then calculates the similarity between the query $q$ and the item $x$ by applying adaptive gating weights, $\pi_{p}(q,x)\in[0,1]$, to the inner products of these $P$ pairs of low-rank embeddings, or $\langle f_{p}(q),g_{p}(x)\rangle$s.
+Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ [42, 3], and parameterizes $\pi_{p}(q,x)$ utilizing some neural network with $f_{p}(q)$s, $g_{p}(x)$s, and $\langle f_{p}(q),g_{p}(x)\rangle$s as input features, but these details do not affect our analyses in this paper.
+Following [42]:
 $\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ (1)
-To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings (where $P_{q}\times P_{x}=P$) and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
+To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embedding, where $P_{q}\times P_{x}=P$, and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
 $\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$ (2)
-We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper as the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
+We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper, given that the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
 Figure 1: Mixture-of-logits (MoL) learned similarity.
-Now we show that any high rank matrix can be decomposed into a mixture of logits with low rank matrices. Specifically, without loss of generality, we prove the following:
+Now we show that any high rank matrix can be decomposed into a mixture of logits based on low rank matrices. Specifically, without loss of generality, we prove the following:
 Theorem 1.
 […]
-First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$:
+We can think about $n$ as the number of queries and $m$ the number of items (or vice versa). First, the theorem trivially holds if the rank of $A$ is less than or equal to $d$ ($d\leq n$):
 Lemma 1.
@@
 Let $\lambda_{min}=\min{\lambda_{ij}}$, and $\lambda_{max}=\max{\lambda_{ij}}$.
-Let $B_{1}=\lambda_{min}UI_{d}V$, $B_{2}=\lambda_{max}UI_{d}V$, where $I_{d}$ is the identity matrix of rank $d$.
+Let $B_{1}=\lambda_{min}U{D_{n,d}}V$, $B_{2}=\lambda_{max}U{D_{n,d}}V$, where $D_{n,d}$ denotes an $n$-by-$n$ diagonal matrix with the first $d$ elements of the diagonal being $1$s and the rest being 00s.
 We have $A^{\prime}_{ij}=\lambda_{ij}\sum_{k=1}^{d}U_{ik}V_{kj}=\pi(i,j)\cdot B_{1ij}+(1-\pi(i,j))\cdot B_{2ij}$, where $\pi(i,j)=\frac{\lambda_{max}-\lambda_{ij}}{\lambda_{max}-\lambda_{min}}$.
@@
 We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4).
-Then we evaluate the MoL score of the items in the set $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8).
+Then we evaluate the MoL scores of the items in $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8).
 Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11).
-Finally, we evaluate the MoL scores of the items in $G^{\prime}$ exhaustively, and return the top $K$ items with the highest scores (line 12).
+Finally, we evaluate the MoL scores of the items in $G^{\prime}$, and return the top $K$ items with the highest scores (line 12).
 We argue that Algorithm 1 retrieves the exact top $K$ items with MoL.
@@
 Note that $K^{\prime}$ can be larger than $K$, e.g., due to oversampling.
-In the second stage, we evaluate the MoL score of the candidate items and return the top $K$ items (line 3).
+In the second stage, we evaluate the MoL scores of the candidate items and return the top $K$ items (line 3).
 Algorithm 2 Approximate top $K$ algorithms.
@@
 In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3.
-Our code and pretrained model checkpoints are publicly available at https://github.com/bailuding/rails.
 4.1 Setup
 Workloads.
-We consider a classical retrieval task (next-item prediction) in recommendation systems (Covington et al., 2016; Zhu et al., 2018; Li et al., 2019; Zhai et al., 2023).
-We employ three widely used datasets, MovieLens (Harper and Konstan, 2015) (the 1M and 20M subsets), and Amazon Reviews (McAuley et al., 2015) (the largest Books subset).
-We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets (Kang and McAuley, 2018; Zhai et al., 2023).
-In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step.
+We consider a classical retrieval task (next-item prediction) in recommendation systems [5, 45, 26, 42].
+We employ three widely used datasets, the 1M and 20M subsets of MovieLens [13], and the largest Books subset of Amazon Reviews [29].
+We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets [20, 42].
+In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step, which is then mapped to $P_{q}$ embeddings using an multi-layer perceptron (MLP) when using the MoL learned similarity, consistent with prior work.
+We discuss detailed hyperparameter settings in Appendix A.
 Table 2 shows the statistics of these three workloads.
-Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ component-level embedding dimension
+Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ $d_{P}$
 ML-1M 6,040 3,649 8 4 64
@@
 Table 2: Workload statistics.
-Model training.
-We implemented MoL on top of two popular self-attention sequential encoder architectures, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024).
-The models are implemented in PyTorch and are trained with 2x NVIDIA RTX 6000 Ada GPUs.
+Model.
+We implemented MoL on top of two popular sequential architectures, SASRec [20] and HSTU [43], with cosine similarity as the default similarity function (SASRec / HSTU rows in Table 3).
+The models are are implemented in PyTorch and trained with 2x NVIDIA RTX 6000 Ada GPUs.
 Method HR@1 HR@10 HR@50 HR@200 MRR
-ML-1M SASRec .0610 .2806 .5437 .7498 .1307
-SASRec + MoL .0672 .3019 .5711 .7650 .1431
-HSTU .0774 .3306 .5897 .7832 .1571
-HSTU + MoL .0875 .3412 .6013 .7877 .1692
-ML-20M SASRec .0674 .2889 .5503 .7661 .1388
-SASRec + MoL .0731 .3068 .5645 .7768 .1498
-HSTU .0965 .3556 .6143 .8074 .1800
-HSTU + MoL .1014 .3661 .6234 .8116 .1876
+ML-1M SASRec .0610 .2818 .5470 .7540 .1352
+SASRec + MoL .0710 .3034 .5655 .7680 .1448
+HSTU .0750 .3332 .5956 .7824 .1579
+HSTU + MoL .0837 .3460 .6048 .7874 .1673
+ML-20M SASRec .0653 .2883 .5484 .7658 .1375
+SASRec + MoL .0790 .3124 .5685 .7784 .1551
+HSTU .0962 .3557 .6146 .8080 .1800
+HSTU + MoL .1008 .3697 .6244 .8124 .1880
 Books SASRec .0058 .0306 .0754 .1431 .0153
@@
 It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3).
-Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 21.4% in HR@1, an average of 13.7% in HR@10, and an average of 16.2% in MRR across the six scenarios, validating the importance of learned similarity functions.
-We evaluate our proposed retrieval algorithms relative to the best performing rows, specifically “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the following section.
+Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 24.2% in HR@1, 14.3% in HR@10, and 16.3% in MRR across the six scenarios, validating that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable.
+We evaluate our proposed retrieval algorithms relative to the best performing rows, “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the next section.
 4.2 Top $K$ retrieval performance
@@
 The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$.
-We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
+We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms [19, 4, 42].
 We set batch size to $32$ for ML-1M, ML-20M, and Books.
@@
 Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ can be used if the target HR is extremely stringent.
-Limitations
-[…]
 5 Related work
-Retrieval in Recommendation Models.
-Most modern recommendation models follow a classical two-stage information retrieval paradigm (Covington et al., 2016), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019) are first filtered down to hundreds to thousands in the retrieval or candidate generation stage, followed by the ranking stage that produces the final top 1-10 candidates.
-Earlier work on scaling large-scale recommendation models primarily utilize dual-encoder setups(Covington et al., 2016) or sometimes graph algorithms (Eksombatchai et al., 2018).
-Researchers quickly realized that dot products limited retrieval stage’s performance, and introduced multi-query embeddings (Li et al., 2019) and specialized neural networks (He et al., 2017) to alleviate the expressiveness constraints in retrieval.
-Learned index structures, or generative retrieval, map each candidate item to a series of tokens and leverage beam search to replace approximate nearest neighbor search.
-These tokens are represented either implicitly during tree traversal (Zhu et al., 2018; Zhuo et al., 2020) or explicitly (Gao et al., 2021).
-However, they introduce stability challenges as recommendation systems need to support billion-scale dynamic realtime updated set of items.
-A latest trend is to exploit characteristics of GPUs, TPUs, and other accelerators; learned similarity functions as well as hardware-aware algorithms have been shown to achieve significant quality improvements (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
+Similarity Functions in Retrieval.
+Most information retrieval models in recommendations and natural language processing follow a classical two-stage paradigm [5], where up to billions of items [6, 26, 3] are first filtered down to hundreds in the retrieval stage, followed by the ranking stage that produces the final top 1-10 candidates.
+Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function [5, 21].
+Researchers quickly realized that dot products limited retrieval stage’s performance, and explored various learned similarity-based approaches.
+Prominent variants include maximum similarity based on multiple embeddings [26, 22, 34], specialized neural networks, often based on Hadamard products [14, 37, 39, 3], and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal [16, 45, 46] or explicitly in the “generative retrieval” setups [8, 38, 40, 36].
+It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendations [33] and DSI in NLP [36].
+Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items [3].
+Despite these challenges, significant gains (17% gains at Hit Rate@100 [42] to 24% gains at Hit Rate@400 [3]) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions [34, 42], implicit diversification done as part of beam search [8, 36], explicit representation of side-information using special neural architectures [3], and hardware-aware similarity function and inference algorithm design on GPUs [4, 42, 3].
 Efficient Nearest Neighbor Search (NNS).
@@
 Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS).
-Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on.
-The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees while maintaining max-heap property have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020).
-Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize GPUs; to address this issue, accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024).
+Various techniques were proposed and analyzed, including tree structures [1, 32], locality sensitive hashing [10, 35], production quantization [18, 11], data partitioning [27, 44], graph-based methods [28, 17], and so on.
+The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results [46].
+Algorithms based on implicit [28, 17, 39, 31] or explicit graphs [39] have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case [30].
+Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques [35, 11] no longer fully utilize GPUs; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work [4, 42, 31, 3].
 6 Conclusion
 We have analyzed techniques for efficient retrieval with learned similarities in this work.
-We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and empirically MoL consistently outperforms dot products in HR across 6 scenarios by an average of 21.4% in HR@1 and 13.7% in HR@10, setting new state-of-the-art on recommendation datasets.
+We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and further empirically learnable – MoL consistently outperforms dot products in HR across 6 scenarios by an average of 24.2% in HR@1 and 14.3% in HR@10, setting new state-of-the-art on recommendation datasets.
 We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds.
@@
 We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations.
-We believe this result is of interest to both recommendation system research communities and researchers working on enabling advanced learned similarities in vector databases.
+Given MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items [42, 3], our work concludes both theoretical and practical justifications for migrating vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.
+Limitations
+[…]
+Appendix A Experiment Setups
+Reproducibility.
+Our code is publicly available online, and detailed implementations and hyperparameter settings for reproducing our experiment results can be found at the following url: https://github.com/bailuding/rails.
+We discuss key settings below.
+Parameterization of low-rank (“component-level”) embeddings, $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$.
+We note that embedding parameterization was not explicitly discussed in the original Mixture-of-Logits paper [42].
+A specific implementation utilizing cluster id information has been proposed and discussed in a follow up work [3], where in the context of large-scale recommendation systems, the authors suggested cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$.
+However, we cannot easily access similar information in the publicly available MovieLens and Amazon Reviews datasets.
+We therefore follow implementation provided by the authors [42] and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems [24].
+All other $f_{p}(q)$s and $g_{p}(x)$s are obtained by applying an MLP on top of query-side/item-side representations in standard sequential recommendation setups [15, 20].
+Parameterization of $\pi_{p}(q,x)$ matrices.
+We follow the implementation provided by the authors [42], which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU [7] non-linearity.
+The inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the learned low-rank embeddings.
+Hyperparameter settings.
+We use an identical number of sampled negatives for dot product baselines (cosine similarity, “SASRec”, “HSTU” rows in Table 3) and Mixture-of-logits (“SASRec + MoL”, “HSTU + MoL” rows in Table 3) to ensure a fair comparison, which is $128$ for ML-1M and ML-20M and $512$ for Amazon Books following prior work.
+For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$), whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets.
+Due to resource constraints, we followed initial hyperparameters provided by the authors [42] for all other parameters.
+A.1 Efficiency
+Method HR@10 HR@50 MRR Time/Epoch
+ML-1M SASRec .2818 .5470 .1352 37s
+SASRec + MoL .3034 (+7.7%) .5655 (+3.4%) .1448 (+7.1%) 44s (+18.9%)
+HSTU .3332 .5956 .1579 41s
+HSTU + MoL .3460 (+3.8%) .6048 (+1.5%) .1673 (+6.0%) 46s (+12.2%)
+ML-20M SASRec .2883 .5484 .1375 197s
+SASRec + MoL .3124 (+8.4%) .5685 (+3.7%) .1551 (+12.8%) 449s (+127.9%)
+HSTU .3557 .6146 .1800 526s
+HSTU + MoL .3697 (+3.9%) .6244 (+1.6%) .1880 (+4.4%) 773s (+47.0%)
+Books SASRec .0306 .0754 .0153 506s
+SASRec + MoL .0402 (+31.4%) .0970 (+28.6%) .0203 (+32.7%) 964s (+90.5%)
+HSTU .0469 .1066 .0233 1883s
+HSTU + MoL .0613 (+30.7%) .1292 (+21.2%) .0315 (+35.2%) 2236s (+18.7%)
+Table 5: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
+Note that these results are identical to those reported in Table 3 but with relative gains (MoL vs dot product) and training cost changes illustrated under different sequential encoder backbones.
+Training efficiency.
+We measure the training efficiency of the setups in Section 4 over one epoch of data, utilizing a single RTX 6000 Ada GPU and with a constant batch size of 128 for ML-1M and ML-20M, and 64 for Amazon Books in Table 5.
+Consistent with prior findings [42, 3], MoL’s model quality increases do not generally come with a large increase in training time, thanks to MoL’s higher arithmetic intensity on GPUs, with an average increase of 53% across the six scenarios.
+To contextualize this 1.53x training cost, prior work on sequential recommenders, such as self-supervised bidirectional models, came with a 5.9x training time increase on ML-20M and 4.5x training time increase on Amazon Reviews (BERT4Rec vs SASRec rows) [23].
+We expect additional GPU kernel-level optimizations, such as kernel fusions, to further improve efficiency of the “+ MoL” rows.
+Inference efficiency.
+Given the number of items (3K, 24K, and 674K) in these datasets, as discussed in Table 2, these public benchmark datasets are not good candidates for measuring inference latency.
+We note that prior work have reported batched inference latency between 5ms and 20ms on datasets with hundreds of millions to billions of items [42, 3], which is inline with optimized MIPS baselines.

### v4

@@
-Efficient Retrieval with Learned Similarities
-Abstract
-Retrieval plays a fundamental role in recommendation systems, search, and natural language processing by efficiently finding relevant items from a large corpus given a query.
-Dot products have been widely used as the similarity function in such retrieval tasks, thanks to Maximum Inner Product Search (MIPS) that enabled efficient retrieval based on dot products.
+Retrieval with Learned Similarities
+Abstract.
+Retrieval plays a fundamental role in recommendation systems, search, and natural language processing (NLP) by efficiently finding relevant items from a large corpus given a query.
+Dot products have been widely used as the similarity function in such tasks, enabled by Maximum Inner Product Search (MIPS) algorithms for efficient retrieval.
 However, state-of-the-art retrieval algorithms have migrated to learned similarities.
-Such algorithms vary in form; the queries can be represented with multiple embeddings, complex neural networks can be deployed, the item ids can be decoded directly from queries using beam search, and multiple approaches can be combined in hybrid solutions.
+These advanced approaches encompass multiple query embeddings, complex neural networks, direct item ID decoding via beam search, and hybrid solutions.
 Unfortunately, we lack efficient solutions for retrieval in these state-of-the-art setups.
-Our work investigates techniques for approximate nearest neighbor search with learned similarity functions.
-We first prove that Mixture-of-Logits (MoL) is a universal approximator, and can express all learned similarity functions.
-We next propose techniques to retrieve the approximate top $K$ results using MoL with a tight bound.
-We finally compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
-1 Introduction
+Our work addresses this gap by investigating efficient retrieval techniques with expressive learned similarity functions.
+We establish Mixture-of-Logits (MoL) as a universal approximator of similarity functions, demonstrate that MoL’s expressiveness can be realized empirically to achieve superior performance on diverse retrieval scenarios, and propose techniques to retrieve the approximate top-$k$ results using MoL with tight error bounds.
+Through extensive experimentation, we show that MoL, enhanced by our proposed mutual information-based load balancing loss, sets new state-of-the-art results across heterogeneous scenarios, including sequential retrieval models in recommendation systems and finetuning language models for question answering; and our approximate top-$k$ algorithms outperform baselines by up to $\times6$ in latency while achieving >.$>.99$ recall rate compared to exact algorithms.
+[^1] Our code and model checkpoints are available at https://github.com/bailuding/rails.
+Nearest Neighbor Search, Learned Similarities, Top-K Retrieval, Vector Databases, Recommendation Systems, Question Answering
+copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714822isbn: 979-8-4007-1274-6/25/04ccs: Information systems Similarity measuresccs: Information systems Top-k retrieval in databasesccs: Information systems Learning to rankccs: Information systems Probabilistic retrieval modelsccs: Information systems Question answeringccs: Information systems Recommender systemsccs: Computing methodologies Natural language processingccs: Information systems Personalization
+1. Introduction
 Retrieval requires efficient storing, indexing, and querying relevant candidate items represented by high-dimensional vectors.
-Retrieval is widely used as the initial preprocessing stage for applications such as recommendations and natural language processing that operate over corpus with up to billions of items [5, 9, 26, 25, 2].
-In many concrete use cases, such as vector databases [19], the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
-Despite the popularity of dot products and numerous work done to improve their efficiency [18, 35, 28, 4], state-of-the-art retrieval algorithms have long moved to various learned similarity functions.
-Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values [26, 22].
-As another example, Probabilistic Label Trees (PLTs) [16] and Tree-based Deep Models (TDMs) [45, 46] map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf.
-More recent work on generative retrieval in recommendations and natural language processing directly map the query to the item ids in sequence-to-sequence or decoder-only setups [38, 40].
-Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products [8].
-Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms [14, 41, 34, 42].
+Retrieval is widely used as the initial preprocessing stage for internet applications such as recommendations, search, question answering, and natural language processing that operate over corpus with up to billions of items (Covington et al., 2016; Gillick et al., 2018; Li et al., 2019; Karpukhin et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022).
+In many concrete use cases, such as vector databases (Johnson et al., 2021), the query- and the item- embeddings are learned with deep neural networks in a dual-encoder setup, and dot products are applied on top of such embeddings as the similarity function for measuring relevance.
+Despite the popularity of dot products and numerous work done to improve their efficiency (Jegou et al., 2011; Shrivastava and Li, 2014; Malkov and Yashunin, 2020; Chern et al., 2022), state-of-the-art retrieval algorithms have long moved to various learned similarity functions.
+Their most basic versions preserve some dot product-related structures, but turn either the query or the item into multiple embeddings, and rely on a max operator to combine those similarity values (Li et al., 2019; Khattab and Zaharia, 2020).
+As another example, Probabilistic Label Trees (PLTs) (Jasinska et al., 2016) and Tree-based Deep Models (TDMs) (Zhu et al., 2018; Zhuo et al., 2020) map items to leaf nodes in a tree, and reduce retrieval to beam search by making decisions sequentially using learned classifiers while traversing trees from root to leaf.
+More recent work on generative retrieval directly map the query to the item ids in sequence-to-sequence or decoder-only setups (De Cao et al., 2021; Bevilacqua et al., 2022; Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023).
+Combinations of these approaches have also been studied, with some performing coarse-grained retrieval with generative approaches, followed by re-ranking using dot products (Gao et al., 2021).
+Finally, the similarity function can be directly parameterized by carefully designed deep neural networks that take various forms (He et al., 2017; Yang et al., 2018; Santhanam et al., 2022; Zhai et al., 2023).
 Supporting efficient retrieval with these diverse learned similarities is challenging.
-Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps [45, 42], while recommendation and NLP deployments commonly need to handle billions of items [6, 26, 3] with low end-to-end latency.
-When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force [14] or heuristics [42].
-What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult for designing a general solution for efficient retrieval.
-Taking a step back, our key insight is that learned similarity approaches are but different ways to increase expressiveness of the retrieval stage.
+Learned similarity functions are generally expensive to compute; with learned index structures, traversing a binary tree with 4 million items requires running beam search for 20 non-parallelizable steps (Zhu et al., 2018), while recommendation and NLP deployments commonly need to handle billions of items (Eksombatchai et al., 2018; Li et al., 2019; Borisyuk et al., 2024) with a latency budget of tens of milliseconds.
+When an arbitrary deep neural network is employed, it’s no longer clear how to perform top-$K$ retrieval other than through brute-force (He et al., 2017) or heuristics (Zhai et al., 2023).
+While graph-based methods can be used to prune the search space (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019; Wang et al., 2022a; Ootomo et al., 2024), such methods tend to be much slower compared with MIPS algorithms leveraging quantization at high recall rates (Guo et al., 2020; ann, [n.
+d.]), and their performance can degrade when the similarity function is not a distance metric (Morozov and Babenko, 2018).
+What is worse, these algorithms vary significantly in terms of their exact formulations, and the lack of a universal interface makes it even more difficult to design a general solution for efficient retrieval.
+Taking a step back, our key insight is that learned similarity approaches are but different ways to increase the expressiveness of the retrieval stage.
 Formally, for a query $q$ and an item $x$, the expressiveness of the similarity function boils down to deriving alternative parameterizations of $p(x|q)$ matrices, with full rank matrices being the most expressive among them.
 Dot products, on the other hand, induces a low-rank bottleneck due to the dimensionality of the embedding, i.e., $\ln p(x|q)\propto\langle f(q),g(x)\rangle$ ($f(q),g(x)\in\mathbb{R}^{d}$).
-This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases [19, 4], and overfitting issues that come with larger embedding dimensions due to the need to co-train query- and item- encoders from data [5, 26, 8, 43].
-This insight enables us to support efficient retrieval with learned similarity functions by approximating them with MoL.
-To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities.
-We first show that Mixture-of-Logits (MoL) is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions.
-Compared with prior work on learned similarities [14, 38] which often fail to outperform dot product baselines in large-scale settings [33, 36], our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on corpus with hundreds of millions to billions of items [42, 3].
-We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight bound.
-Our solution leverages existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS [18, 35, 19, 12].
-We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to 91×91$\times$91 × in latency, while achieving >.$>.99$ recall rate of exact algorithms.
-Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity [42], which results in MIPS-level inference latency and throughput.
-This concludes both theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
-2 Expressiveness of Mixture of Logits
-In this section, we describe Mixture of Logits (MoL) and show that MoL is expressive enough to represent any learned similarity function. Table 1 summarizes the notations in this paper.
-We first describe the learned similarity function with Mixture of Logits (MoL).
+This cannot be alleviated by simply increasing the embedding dimension $d$, due to memory bandwidth being the main bottleneck in modern dot-product based retrieval systems, such as vector databases (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023), and overfitting issues that come with larger embedding dimensions due to the common need to co-train or finetune query- and item- encoders from data (Covington et al., 2016; Li et al., 2019; Karpukhin et al., 2020; Gao et al., 2021; Ni et al., 2022a, b; Zhai et al., 2024).
+This insight enables us to support efficient retrieval with expressive learned similarity functions by approximating them with Mixture-of-Logits (MoL).
+To the best of our knowledge, this is the first work that tackles the problem of efficient retrieval with universal learned similarities, while setting new state-of-the-art results across heterogeneous scenarios.
+We first show that Mixture-of-Logits is a universal approximator as it can express $p(x|q)$ matrices of arbitrary high rank, and hence approximate all learned similarity functions (Section 2.1).
+Our work lays theoretical foundations for MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 on web-scale corpus with hundreds of millions to billions of items (Zhai et al., 2023; Borisyuk et al., 2024), and further enables MoL to be effectively applied across diverse retrieval scenarios, from large-scale recommendation systems to finetuning language models for question answering (Section 2.2).
+We next propose techniques to retrieve the approximate top-$K$ results using MoL with a tight error bound (Section 3).
+Our solution leverages the existing widely used APIs of vector databases like top-K queries, thus benefiting from prior work on efficient vector search like MIPS (Jegou et al., 2011; Shrivastava and Li, 2014; Johnson et al., 2021; Guo et al., 2020).
+We empirically compare our techniques with existing approaches, showing that MoL sets new state-of-the-art results on recommendation retrieval and question answering tasks, and our approximate top-k retrieval with learned similarities outperforms baselines by up to $\times6$ in latency, while achieving >.$>.99$ recall rate of exact algorithms (Section 4).
+Importantly, our approach with learned similarities efficiently utilizes modern accelerators due to MoL’s higher arithmetic intensity (Zhai et al., 2023), which results in MIPS-level inference latency and throughput.
+Overall, our work provides strong theoretical and practical justifications to migrate away from the broadly adopted MIPS solution in vector databases to Retrieval with Learned Similarities (RAILS) on GPUs.
+2. Mixture of Logits
 Notation Description
@@
 $\phi(q,x)$ the learned similarity function, i.e., Mixture-of-Logits (MoL).
-$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ low-rank embeddings ("component-level embeddings") to represent $q$ and $x$.
-With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively, and $P=P_{q}\times P_{x}$.
+$P$ ($P_{q}$, $P_{x}$) MoL uses $P$ pairs of low-rank embeddings ("component-level embeddings") to represent $q$ and $x$.
+With the (batched) outer product form of MoL, $P_{q}$ and $P_{x}$ are the numbers of embeddings for $q$ and $x$, respectively; $P=P_{q}\times P_{x}$.
 $\pi_{p}(q,x)$ ($\pi_{p_{q},p_{x}}(q,x)$) weight for the $p$-th (or $p_{q}$-th by $p_{x}$-th with outer product) embedding set for $(q,x)$.
@@
 $d_{P}$ dimensionality of low-rank (component-level) embeddings. $f_{p}(q),g_{p}(q)\in\mathbb{R}^{d_{P}}$.
-$\langle f(q),g(x)\rangle$ the dot product similarity function; $\langle f(q),g(x)\rangle=g(x)^{T}f(q)$.
-Table 1: Table of Notations.
+$\langle f(q),g(x)\rangle$ the dot product similarity function: $g(x)^{T}f(q)$. $\langle f_{p}(q),g_{p}(x)\rangle$ denotes the dot product for the $p^{th}$ embedding pair.
+Table 1. Table of Notations.
+In this section, we describe Mixture of Logits (MoL), propose a load balancing loss to improve conditional computations in MoL, prove that MoL is expressive enough to represent any learned similarity function, and demonstrate how to apply MoL to diverse retrieval tasks.
+Table 1 summarizes the notations in this paper.
+We first describe Mixture of Logits (MoL).
 Mixture of Logits (MoL)
-MoL [42, 3] assumes that the query $q$ and the item $x$ are already mapped to $P$ groups of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively.
+MoL (Zhai et al., 2023) assumes that the query $q$ and the item $x$ are already mapped to $P$ pairs of low-rank embeddings (“component-level embeddings”), $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, where $f_{p}(q),g_{p}(x)$ are parameterized with some neural networks based on query and item features, respectively, and $d_{P}$ is the dimensionality of the low-rank embeddings.
 MoL then calculates the similarity between the query $q$ and the item $x$ by applying adaptive gating weights, $\pi_{p}(q,x)\in[0,1]$, to the inner products of these $P$ pairs of low-rank embeddings, or $\langle f_{p}(q),g_{p}(x)\rangle$s.
-Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ [42, 3], and parameterizes $\pi_{p}(q,x)$ utilizing some neural network with $f_{p}(q)$s, $g_{p}(x)$s, and $\langle f_{p}(q),g_{p}(x)\rangle$s as input features, but these details do not affect our analyses in this paper.
-Following [42]:
-$\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$ (1)
-To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embedding, where $P_{q}\times P_{x}=P$, and applying l2-norm to $f_{p}(q)$s and $g_{p}(x)$s:
-$\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$ (2)
+Note that prior work assumes that $\sum_{p}\pi_{p}(q,x)=1$ (Zhai et al., 2023; Borisyuk et al., 2024), but this does not affect our analyses in this paper.
+Following (Zhai et al., 2023):
+(1) $\phi(q,x)=\sum_{p=1}^{P}\pi_{p}(q,x)\langle f_{p}(q),g_{p}(x)\rangle$
+To extend this to large-scale datasets and to enable hardware-efficient implementations on accelerators like GPUs, Equation 1 was further modified by decomposing those $P$ dot products as (batched) outer products of $P_{q}$ query-side and $P_{x}$ item-side embeddings, where $P_{q}\times P_{x}=P$, and applying l2-norm to the embeddings:
+(2) $\phi(q,x)=\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}\pi_{p_{q},p_{x}}(q,x)\left\langle\frac{f_{p_{q}}(q)}{||f_{p_{q}}(q)||_{2}},\frac{g_{p_{x}}(x)}{||g_{p_{x}}(x)||_{2}}\right\rangle$
 We use Equation 1 and 2 interchangeably as the MoL form to analyze throughout the rest of this paper, given that the embedding normalization for $f_{p_{q}}(q)$s and $g_{p_{x}}(x)$s can be precomputed.
-Figure 1: Mixture-of-logits (MoL) learned similarity.
-Now we show that any high rank matrix can be decomposed into a mixture of logits based on low rank matrices. Specifically, without loss of generality, we prove the following:
+Mixture of Logits (MoL) with load balancing regularization loss.
+We further observe $\pi_{p}(q,x)$ defines conditional computation to be performed over the $p$ low-rank embedding pairs, or $(f_{p}(q),g_{p}(x))$s. $\pi_{p}(q,x)$ should hence satisfy two conditions:
+  • Globally, the $p$ low-rank embedding pairs, or $(f_{p}(q),g_{p}(x))$s, should receive a similar number of training examples even when $p$ is large and $\pi_{p}(q,x)$ is sparse, with load distributed evenly across the $p$ pairs.
+One way to do this is to maximize the entropy $H(p)$ over these embedding pairs.
+  • The low-rank embedding pairs used to compute a particular $\phi(q,x)$ should be non-uniform and ideally sparse; e.g., it’s desirable to avoid the degenerate solution where $\pi_{p}(q,x)=\frac{1}{p}$.
+One way to do this is to minimize the conditional entropy $H(p|(q,x))$ of $p$ given (query, item) pairs.
+Given these two desired conditions, we propose a mutual information-based regularization loss for load balancing, defined as
+(3) $\mathcal{L}_{MI}=-H(p)+H(p|(q,x))$
+with the overall training loss as
+(4) $-\log\frac{\exp(\phi(q,x))}{\exp(\phi(q,x))+\sum_{x^{\prime}\in\mathbb{X}}\exp(\phi(q,x^{\prime}))}+\alpha\mathcal{L}_{MI}$
+where the first part of Equation 4 is the sampled softmax loss used in (Zhai et al., 2023), and the second part adjusts the weight for the mutual information-based load balancing loss with a hyperparameter $\alpha$.
+Figure 1. Mixture-of-logits (MoL) learned similarity.
+2.1. Expressiveness of Mixture of Logits
+Now we show that any high-rank matrix can be decomposed into a mixture of logits based on low-rank matrices, i.e., MoL is a universal approximator for all similarity functions.
+Without loss of generality, we prove the following:
 Theorem 1.
@@
 Note that our decomposition is not intended to be used as a distillation of the original high-rank matrix.
-For instance, it is prohibitively expensive to populate the full matrix with a learned similarity function.
-In addition, the proof also does not indicate that having a sum of two mixture of logits is sufficient to train the embeddings and the learned similarity function.
+It is likely prohibitively expensive to populate the full matrix with a learned similarity function.
+In addition, our proof also does not indicate that having two mixture components is sufficient to train the embeddings and the learned similarity function.
 It is well-known that overparameterization is often necessary to enable efficient and performant training.
-3 Retrieval Algorithms
-In this section, we describe the problem of retrieving the top $K$ items with MoL and both exact and approximate algorithms for retrieval. Formally, we define the top $K$ retrieval problem as follows:
+2.2. Applying MoL to Heterogeneous Use Cases
+We now discuss how to apply MoL to retrieval tasks in different domains.
+Parameterization of the low-rank, component-level embeddings, or $f_{p}(q),g_{p}(x)\in\mathbb{R}^{d_{P}}$, plays an important role in realizing MoL’s theoretical expressiveness in practice, as suggested by prior work (Borisyuk et al., 2024).
+We discuss two scenarios on the opposite end of the spectrum, one with a large number of heterogeneous features – retrieval in large-scale recommendation systems, followed by another with a single homogeneous feature – finetuning language models for question answering and related NLP use cases, shown in Figure 2.
+Retrieval in Large-scale Recommendation Systems.
+Recommendation systems are characterized by the large number of heterogeneous features they use (Covington et al., 2016; Song et al., 2019; Zhai et al., 2024).
+This naturally enables some of those features to be utilized on the query- (user-) or on the item-side.
+For instance, embeddings can be constructed based on cluster ids on both the query-side and the item-side (Borisyuk et al., 2024).
+For common benchmark datasets, User ID-based one-hot embeddings (Koren et al., 2009) represent another possible $g_{p}(q)$ to use, which we evaluate in Section 4.
+Finetuning Language Models for Question Answering.
+In contrast, language models are characterized by their use of homogeneous semantic features, such as wordpieces and sentencepieces (Kudo and Richardson, 2018).
+We observe that MoL can be similarly adopted for those use cases.
+To obtain the $P_{X}$ item embeddings for MoL, we expand the tokenizer’s vocabulary with $P_{X}$ special aggregation tokens $X_{1},\ldots,X_{P_{X}}$, and append those $P_{X}$ tokens at the beginning of every tokenized sequence, $SP_{1},\ldots,SP_{N}$, as illustrated in Figure 2  [^1] Note that many question answering scenarios (Karpukhin et al., 2020; De Cao et al., 2021; Wang et al., 2022b; Ni et al., 2022b; Sun et al., 2023) utilize bidirectional language models for retrieval, like BERT (Devlin et al., 2019) or T5 (Raffel et al., 2023); for recent unidirectional language models, we can add $X_{1},\ldots,X_{P_{X}}$ to the end of the input sequence instead..
+These $P_{X}$ special tokens play similar roles as the CLS token in BERT (Devlin et al., 2019), and during finetuning of the language model, are co-trained to aggregate different aspects of information as inputs for MoL.
+Additionally, we can design a learned pooling function to adapt pooling policy at an example-level (“Parameterized Pooling”) to improve model quality, which we discuss further in Appendix A.2.
+Figure 2.
+Illustration of how to apply Mixture-of-logits (MoL) learned similarity to various retrieval scenarios, with a language model finetuning use case (characterized by a single homogeneous feature) shown on the left, and a recommendation use case (characterized by a large number of heterogeneous features) shown on the right.
+More details can be found in Appendix A.2.
+3. Retrieval Algorithms
+In this section, we describe the problem of retrieving the top $K$ items with MoL as well as exact and approximate retrieval algorithms. Formally, we define the top $K$ retrieval problem as follows:
 Definition 0.
@@
 Gap of approximate top $K$: Let $q$ be a query and $X_{K}$ be the set of exact top $K$ items for the query $q$ from a set of items $X$.
-Let $X^{*}$ be the approximate top $K$ results, where $X^{*}\subseteq X$.
+Let $X^{*}$ be the set of approximate top $K$ results, where $X^{*}\subseteq X$.
 Let $S=\min\{\phi(q,x),x\in X^{*}\}$ and $S^{\prime}=\max\{\phi(q,x),x\in X_{K}\setminus X^{*}\}$.
 We call $S_{\Delta}=S^{\prime}-S$ the gap of the top $K$ with $X^{*}$.
-3.1 Exact algorithm
+3.1. Exact algorithm
 The brute-force algorithm to retrieve the exact top $K$ with MoL is to evaluate $\phi(q,x)$ for each query $q$ and item $x$.
 This algorithm can be prohibitively expensive if the number of items is large.
-Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1.
+Instead, we describe a more efficient two-pass algorithm to retrieve the exact top $K$ items as shown in Algorithm 1 (example in Appendix B).
 Algorithm 1 Exact top $K$ algorithm.
@@
 4:for $p\in P$ do
-5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
-6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings
+5:    $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed.
+6:    $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p})$triangleright Retrieve top $K$ items for each pair of embeddings.
 7:$S_{min}\leftarrow\infty$
@@
 12:for $p\in P$ do
-13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items with a dot product higher than or equal to $S_{min}$ for $f_{p}(q)$ from $X_{p}$
-14:return $BruteForceTopKMoL(q,G^{\prime})$$\triangleright$ Retrieve the top $K$ items from $G^{\prime}$ with MoL
+13:    $G^{\prime}\leftarrow G^{\prime}\cup RangeDotProduct(f_{p}(q),S_{min},X_{p})$$\triangleright$ Retrieve all items $x\in X_{P}$ with $\langle f_{p}(q),x\rangle\geq S_{min}$.
+14:return $BruteForceTopKMoL(q,G^{\prime})$triangleright Retrieve the top $K$ items from $G^{\prime}$ with MoL.
 We start by retrieving the top $K$ items with the highest dot product scores for each group of embeddings as the initial candidate set $G$ (line 1-4).
-Then we evaluate the MoL scores of the items in $G$ and find the minimal learned similarity score $S_{min}$ (line 5-8).
+Then we evaluate the MoL scores of the items in $G$ and find the minimal MoL score $S_{min}$ (line 5-8).
 Next we retrieve all items within a distance of $S_{min}$ with the query $q$ as the candidate set $G^{\prime}$ (line 9-11).
@@
 So we have shown that $X_{K}=X^{\prime}$.
-3.2 Approximate algorithms
+3.2. Approximate algorithms
 In the exact algorithm shown in Algorithm 1, we need to retrieve all the items with a dot product higher than or equal to a threshold.
-When the threshold is a loose filter of the item set, which may happen when the dot product scores are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large set of candidate items can still be expensive.
-Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact top $K$ algorithm.
+When the threshold is a loose filter of the item set, which may happen when the dot products are skewed, $G^{\prime}$ can be large, and the evaluation of MoL over a large number of candidates can be expensive.
+Here, we describe two heuristics to approximately retrieve the top $K$ items and analyze their gap against the exact algorithm.
 In both heuristics, we perform a two-stage retrieval as shown in Algorithm 2.
@@
 In the second stage, we evaluate the MoL scores of the candidate items and return the top $K$ items (line 3).
-Algorithm 2 Approximate top $K$ algorithms.
+Algorithm 2 Approximate top-$k$ algorithms.
 1:a query $q$, a set of items $X$
@@
 3:function ApproxTopK($q,X,K,K^{\prime}$)
-4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidate items
-5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items with MoL
-6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
-7:union of top $K$ items over $P$ component-level embeddings by dot product
+4:    $G\leftarrow TopKCandidate(q,X,K^{\prime})$triangleright Retrieve the top $K^{\prime}$ candidates.
+5:    return $BruteForceTopKMoL(q,G,K)$$\triangleright$ Retrieve the top $K$ items from $G$ with MoL.
+6:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the $P$ component-level embedding pairs $f_{p}(q),g_{p}(x)$
+7:union of top $K$ items over $P$ embedding pairs by dot product
 8:function TopKPerEmbedding($q,X,K$)
@@
 10:    for $p\in P$ do
-11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed and materialized.
-12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product
-13:    return $G$
-14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embedding $f_{p}(q),g_{p}(x)$
-15:top $K$ items based on the averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
+11:         $X_{p}\leftarrow\{g_{p}(x),x\in X\}$triangleright Can be preprocessed.
+12:         $G\leftarrow G\cup TopKDotProduct(f_{p}(q),X_{p},K)$$\triangleright$ Retrieve the top $K$ items by dot product for the $p$-th embedding pair.
+13:    return $G$triangleright Dedup’ed set of the top $K$ item for each of the $P$ queries
+14:a query $q$, a set of items $X$, $f_{p}(\cdot)$, $g_{p}(\cdot)$ for constructing the component-level embeddings $f_{p}(q),g_{p}(x)$
+15:top $K$ items with averaged dot product, $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$
 16:function TopKAvg($q,X,K$)
 17:    $q^{\prime}\leftarrow\sum_{p=1}^{P}f_{p}(q)$
-18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed and materialized.
+18:    $X^{\prime}\leftarrow\{\sum_{p=1}^{P}g_{p}(x)/P,x\in X\}$triangleright Can be preprocessed.
 19:    return $TopKDotProduct(q^{\prime},X^{\prime},K)$
-Here, we describe two heuristics to retrieve the candidate item set:
+Here, we describe two heuristics to retrieve the candidate items:
 Top $K$ per embedding.
-Given a query $q$ and a set of items $X$, for each embedding set $p$, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$). Return the union across $P$ queries.
-The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding by dot product. We analyze the gap of this heuristic as follows:
+Given a query $q$ and a set of items $X$, for each of the $p$ embedding pairs, retrieve top $K$ items $X_{K,p}$ based on dot product ($\langle f_{p}(q),g_{p}(x)\rangle$).
+Return the union across $P$ queries.
+The top $K$ per embedding heuristic returns the union of the top $K$ items for each embedding pair (group) by dot product. We analyze the gap of this approach as follows:
 Theorem 3.
-Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\phi(q,x),x\in X_{K+1,p}\}$.
-Let $S_{min}$ be the top $K^{th}$ MoL score of the items in $\cup_{p}X_{K,p}$, then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
-Remark Note that there exists an MoL such that $S_{\Delta}=S-S_{min}$, i.e., when $\pi_{p}(q,x)=1$ for $x_{p}=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p}\}$.
+Upper bound of the gap of top $K$ per embedding: Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S=\max\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p},\forall p\}$.
+Let $S_{K}$ be the $K^{th}$ largest MoL score of the items in $\cup_{p}X_{K,p}$, then we have $S_{\Delta}\leq S-S_{K}$.
+Remark This gap (error bound) bounds the maximal difference between the $K^{th}$ largest MoL score from the set of items retrieved by the heuristic and the actual $K^{th}$ largest MoL score.
+In addition, any retrieved item $x$ with $\phi(q,x)\geq S$ belongs to the actual top $K$ items.
+Note that there exists an MoL such that $S_{\Delta}=S-S_{K}$, i.e., when $\pi_{p}(q,x)=1$ for $x,p=\operatorname*{arg\,max}_{x,p}\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X_{K+1,p}\setminus X_{K,p},\forall p\}$.
 Thus, the upper bound of $S_{\Delta}$ is tight.
+In practice, we can calculate a looser bound with the items retrieved by per embedding top $K$, i.e., from $X_{K,p}$.
+We provide an example in Appendix B.
 Top $K$ average.
-Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
+Given a query $q$ and a set of items $X$, return the top $K$ items with the highest average dot product defined as $\sum_{p}\langle f_{p}(q),g_{p}(x)\rangle/P$.
 Note that the top $K$ average heuristic returns the exact top $K$ items when the gating weight distribution in MoL, $\pi$, is uniform.
@@
 This complements the heuristic that retrieves top $K$ per embedding.
-Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding sets $P=P_{q}\times P_{x}$.
-Formally, let $P=P_{q}\cdot P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$.
+Second, in the setup where the set of embedding pairs is constructed as the outer product of the embeddings of a query and those of an item (Equation 2), the average dot product can be efficiently preprocessed and materialized for the items, and the computation of the top $K$ average is then agnostic to the number of embedding pairs.
+Formally, let $P=P_{q}\times P_{x}$ be the number of embedding pairs, where $P_{q}$ is the number of embeddings of a query $q$ and $P_{x}$ is that of an item $x$.
 The average dot product can be computed as
-$\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}=\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}=\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$ (3)
+(5) $\frac{1}{P}\cdot\sum_{p=1}^{P}{\langle f_{p}(q),g_{p}(x)\rangle}$ $=$ $\frac{1}{P}\cdot\sum_{p_{q}=1}^{P_{q}}\sum_{p_{x}=1}^{P_{x}}{\langle f_{p_{q}}(q),g_{p_{x}}(x)\rangle}$
+(6) $=$ $\frac{1}{P}\cdot\left\langle\sum_{p_{q}=1}^{P_{q}}f_{p_{q}}(q),\sum_{p_{x}=1}^{P_{x}}g_{p_{x}}(x)\right\rangle$
 Thus, we can preprocess the embeddings of the items and the query, so the number of embeddings accessed is $1$ per item for a given query, regardless of the overall number of component-level embeddings used by MoL, i.e., $P$.
@@
 Upper bound of the gap of combined top $K$.
-Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{min}$ as defined in Theorem 2.
+Let $X_{K,p}$ be the top $K$ items of the embedding set $p$ and $S_{K}$ as defined in Theorem 3.
 Let $X^{\prime}_{K}$ be the top $K$ items from top $K$ average.
-Let $S^{\prime}=\max\{\phi(q,x),x\in X\setminus(\cup_{p}X_{K,p}\cup X^{\prime}_{K})\}$.
-Then the gap of $S_{\Delta}\leq S^{\prime}-S_{min}$.
-Remark Similar to Theorem 2, the upper bound of the gap is tight.
+Let $S^{\prime}=\max\{\langle f_{p}(q),g_{p}(x)\rangle,x\in X\setminus(X_{K,p}\cup X^{\prime}_{K}),\forall p\}$.
+Then the gap $S_{\Delta}$ satisfies $S_{\Delta}\leq S^{\prime}-S_{K}$.
+Remark Similar to Theorem 3, the upper bound of this gap is tight.
 In practice, we can configure the $K$ to be different for the two heuristics, i.e., $K_{1}$ and $K_{2}$.
 For example, when the weight distribution $\pi$ is more balanced, $K_{2}$ can be configured to be larger as the top $K$ average approach approximates MoL well while being more computationally efficient.
-4 Evaluation
-In this section, we evaluate the performance of the MoL based learned similarity and the efficiency of our retrieval algorithms discussed in Section 3.
-4.1 Setup
-Workloads.
-We consider a classical retrieval task (next-item prediction) in recommendation systems [5, 45, 26, 42].
-We employ three widely used datasets, the 1M and 20M subsets of MovieLens [13], and the largest Books subset of Amazon Reviews [29].
-We utilize sequential user encoders, as they have been shown to achieve state-of-the-art results on these datasets [20, 42].
-In the next-item prediction setting in sequential recommendations, the query $q$ corresponds to the user representation at a particular time step, which is then mapped to $P_{q}$ embeddings using an multi-layer perceptron (MLP) when using the MoL learned similarity, consistent with prior work.
-We discuss detailed hyperparameter settings in Appendix A.
-Table 2 shows the statistics of these three workloads.
+4. Evaluation
+In this section, we evaluate the performance of MoL based learned similarity with the proposed load balancing loss, and the efficiency of retrieval algorithms discussed in Section 3.
+Our code and model checkpoints are available at https://github.com/bailuding/rails.
+4.1. Workloads
+We benchmark MoL with the proposed load balancing loss $\mathcal{L}_{MI}$, on top of state-of-the-art baselines in recommendation systems and question answering. We describe workloads used below.
+Recommendation Systems.
+We consider three widely used datasets, the 1M and 20M subsets of MovieLens (Harper and Konstan, 2015), and the largest Books subset of Amazon Reviews (McAuley et al., 2015).
+Sequential retrieval models have been shown to achieve state-of-the-art results on these datasets (Hidasi et al., 2016; Kang and McAuley, 2018; Zhai et al., 2024).
+In these settings, sequential encoders, like RNNs or Transformers, are used to map user representations at time $t$ – e.g., in a commonly used setting shown in Figure 2, the list of items in user history up until time $t$, $\Phi_{0},\ldots,\Phi_{t}$ – to $\mathbb{R}^{d}$, and the model is trained to autoregressively predict the next item $x_{t+1}$.
+We hence compare MoL with the proposed regularization loss on top of two popular backbones used for sequential retrieval models, SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024), against cosine similarity baselines.
+We utilize user id-based embeddings discussed in Section 2.2 and MLPs to parameterize the $P_{Q}$ query-side and the $P_{X}$ item-side features.
+Question Answering (QA)
+Natural Questions (NQ) (Kwiatkowski et al., 2019) is commonly used to evaluate state-of-the-art neural retrieval models, including dense retrieval (Karpukhin et al., 2020; Ni et al., 2022b) and generative retrieval (De Cao et al., 2021; Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023) approaches in recent years.
+The most commonly used version (Tay et al., 2022; Wang et al., 2022b; Sun et al., 2023), which we reuse in our work, is often referred to as NQ320k.
+NQ320k consists of 320k query-items pairs, where the items are from Wikipedia pages and the queries are natural language questions.
+We utilize special aggregation tokens discussed in Section 2.2 to parameterize embeddings in MoL, and compare MoL with popular sparse retrieval methods (Robertson and Zaragoza, 2009; Nogueira and Lin, 2019), dense retrieval methods (Karpukhin et al., 2020; Ni et al., 2022a, b), and generative retrieval methods (De Cao et al., 2021; Tay et al., 2022; Bevilacqua et al., 2022; Wang et al., 2022b; Sun et al., 2023).
+Consistent with recent work (Zhuang et al., 2023; Wang et al., 2022b; Sun et al., 2023), we use the pre-trained query generation model from DocT5Query (Nogueira and Lin, 2019) to generate synthetic (query, item) pairs for data augmentation.
+Table 2 summarizes the statistics of these four workloads.
 Workload $|Q|$ $|X|$ $|P_{q}|$ $|P_{x}|$ $d_{P}$
@@
 Books 694,897 674,044 8 8 32
-Table 2: Workload statistics.
-Model.
-We implemented MoL on top of two popular sequential architectures, SASRec [20] and HSTU [43], with cosine similarity as the default similarity function (SASRec / HSTU rows in Table 3).
-The models are are implemented in PyTorch and trained with 2x NVIDIA RTX 6000 Ada GPUs.
-Method HR@1 HR@10 HR@50 HR@200 MRR
-ML-1M SASRec .0610 .2818 .5470 .7540 .1352
-SASRec + MoL .0710 .3034 .5655 .7680 .1448
-HSTU .0750 .3332 .5956 .7824 .1579
-HSTU + MoL .0837 .3460 .6048 .7874 .1673
-ML-20M SASRec .0653 .2883 .5484 .7658 .1375
-SASRec + MoL .0790 .3124 .5685 .7784 .1551
-HSTU .0962 .3557 .6146 .8080 .1800
-HSTU + MoL .1008 .3697 .6244 .8124 .1880
-Books SASRec .0058 .0306 .0754 .1431 .0153
-SASRec + MoL .0083 .0402 .0970 .1813 .0203
-HSTU .0101 .0469 .1066 .1876 .0233
-HSTU + MoL .0150 .0613 .1292 .2167 .0315
-Table 3: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
+NQ320K 307,373 109,739 4 4 768
+Table 2. Workload statistics.
+4.2. Quality of MoL-based Learned Similarity
 Metrics.
 We use Recall (Hit Rate) as the main metric.
-It’s worth noting that HSTU + MoL sets new state-of-the-art results on MovieLens and Amazon Reviews (Table 3).
-Mixture-of-Logits (+ MoL rows) also consistently outperform dot products by an average of 24.2% in HR@1, 14.3% in HR@10, and 16.3% in MRR across the six scenarios, validating that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable.
-We evaluate our proposed retrieval algorithms relative to the best performing rows, “HSTU + MoL”, which exhaustively evaluates $\phi(q,x)$ over all $x$s and corresponds to BruteForce rows in the next section.
-4.2 Top $K$ retrieval performance
+We report Hit Rate@{1, 10, 100} and Mean Reciprocal Rank (MRR) on NQ320K, following (Wang et al., 2022b; Sun et al., 2023), and Hit Rate@{1, 10, 50, 200} on ML-1M, ML-20M, and Books, following (Zhai et al., 2023, 2024).
+Hyperparameter Settings.
+We set the weight $\alpha$ for the proposed load balancing loss $\mathcal{L}_{MI}$ to $0.001$ for all experiments.
+We reuse baseline settings for most other hyperparameters, including learning rate, number of examples used for in-batch negative sampling, etc., with detailed discussions in Appendix A.
+For the NQ320K dataset, we reuse SEAL (Bevilacqua et al., 2022) and NCI (Wang et al., 2022b) results reported by (Wang et al., 2022b), and results for other models as reported by (Sun et al., 2023).
+The Sentence-T5 (Ni et al., 2022a), GENRE (De Cao et al., 2021), DSI (Tay et al., 2022), SEAL (Bevilacqua et al., 2022), DSI+QG (Zhuang et al., 2023), NCI (Wang et al., 2022b), and GenRet (Sun et al., 2023) rows are all finetuned from T5-base, consistent with MoL, to ensure a fair comparison.
+All other results are implemented in PyTorch, and are trained with 1x/2x 48GB GPUs for the recommendation datasets and 4x 80GB GPUs for the QA datasets.
+Method HR@K MRR
+K=1 K=10 K=50 K=200
+ML-1M dataset
+SASRec (Kang and McAuley, 2018) .0610 .2818 .5470 .7540 .1352
+SASRec + MoL .0697 .3036 .5617 .7667 .1441
+HSTU (Zhai et al., 2024) .0750 .3332 .5956 .7824 .1579
+HSTU + MoL .0884 .3465 .6022 .7935 .1712
+HSTU + MoL abl. $\mathcal{L}_{MI}$0847 .3417 .6011 .7942 .1662
+ML-20M dataset
+SASRec (Kang and McAuley, 2018) .0653 .2883 .5484 .7658 .1375
+SASRec + MoL .0778 .3102 .5682 .7779 .1535
+HSTU (Zhai et al., 2024) .0962 .3557 .6146 .8080 .1800
+HSTU + MoL .1010 .3698 .6260 .8132 .1881
+HSTU + MoL abl. $\mathcal{L}_{MI}$0994 .3670 .6241 .8128 .1866
+Books dataset
+SASRec (Kang and McAuley, 2018) .0058 .0306 .0754 .1431 .0153
+SASRec + MoL .0095 .0429 .0915 .1635 .0212
+HSTU (Zhai et al., 2024) .0101 .0469 .1066 .1876 .0233
+HSTU + MoL .0156 .0631 .1308 .2173 .0324
+HSTU + MoL abl. $\mathcal{L}_{MI}$0153 .0625 .1286 .2172 .0321
+Table 3. Evaluation of performance for sequential retrieval models on MovieLens and Amazon Reviews.
+Method HR@K MRR
+K=1 K=10 K=100
+Sparse retrieval
+BM25 (Robertson and Zaragoza, 2009) .297 .603 .821 .402
+DocT5Query (Nogueira and Lin, 2019) .380 .693 .861 .489
+Dense retrieval
+DPR (Karpukhin et al., 2020) .502 .777 .909 .599
+Sentence-T5 (Ni et al., 2022a) .536 .830 .938 .641
+GTR-Base (Ni et al., 2022b) .560 .844 .937 .662
+Generative retrieval
+GENRE (De Cao et al., 2021) .552 .673 .754 .599
+DSI (Tay et al., 2022) .552 .674 .780 .596
+SEAL (Bevilacqua et al., 2022) .570 .800 .914 .655
+DSI+QG (Zhuang et al., 2023) .631 .807 .880 .695
+NCI (Wang et al., 2022b) .659 .852 .924 .731
+GenRet (Sun et al., 2023) .681 .888 .952 .759
+Learned similarities
+MoL .685 .919 .970 .773
+MoL abl. $\mathcal{L}_{MI}$673 .919 .968 .767
+Table 4. Evaluation of performance for QA retrieval models finetuned from language models on Natural Questions.
+Results.
+Across the six recommendation scenarios utilizing different sequential encoder backbones, Mixture-of-Logits (MoL rows) consistently outperform the state-of-the-art dense retrieval baselines (dot products) by an average of 29.1% in HR@1, 16.3% in HR@10, and 18.1% in MRR (Table 3).
+On the widely used Natural Questions QA dataset, MoL outperforms all recent generative retrieval approaches as well as strong dense- and sparse- retrieval baselines (Table 4).
+These results validate that learned similarities, in particular MoL, are not only theoretically expressive but also practically learnable, improving retrieval quality across heterogeneous scenarios, including sequential retrieval models for Recommendations and finetuning LMs for Question Answering.
+Ablation Studies.
+We conduct ablation studies for the proposed mutual information-based load balancing loss relative to the best performing method for each dataset (“abl.
+$\mathcal{L}_{MI}$” rows).
+Results show that our proposed $\mathcal{L}_{MI}$ loss improves HR@1 by 2.4%, HR@10 by 0.8% and MRR by 1.4% across the four datasets.
+In particular, our proposed $\mathcal{L}_{MI}$ loss enables MoL to outperform the best generative retrieval approach on NQ320K, GenRet (Sun et al., 2023), across all metrics.
+4.3. Top $K$ retrieval performance
 Method HR@1 HR@5 HR@10 HR@50 HR@100 Latency / ms
-ML-1M BruteForce 1.00 1.00 1.00 1.00 1.00 .64$\pm$.03
-TopKPerEmbd5 .838 .768 .681 .499 .437 .95$\pm$.03
-TopKPerEmbd10 .932 .905 .842 .678 .618 .94$\pm$.05
-TopKPerEmbd50 1.00 .997 .999 .977 .954 .95$\pm$.05
-TopKPerEmbd100 1.00 1.00 1.00 .995 .990 1.24$\pm$.04
-TopKAvg200 .980 .970 .967 .958 .948 .74$\pm$.06
-TopKAvg500 .992 .986 .989 .990 .990 .71$\pm$.04
-TopKAvg1000 .996 .994 .998 .997 .999 .72$\pm$.03
-CombTopK5_200 1.00 .997 .988 .975 .960 1.09$\pm$.08
-CombTopK50_500 1.00 .999 .999 .998 .997 1.13$\pm$.04
-ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 3.08$\pm$.02
-TopKPerEmbd5 .755 .702 .645 .467 .402 1.19$\pm$.05
-TopKPerEmbd10 .946 .876 .817 .645 .564 1.17$\pm$.05
-TopKPerEmbd50 .993 .986 .977 .931 .900 1.55$\pm$.03
-TopKPerEmbd100 .998 .997 .994 .979 .966 2.26$\pm$.03
-TopKAvg200 .992 .992 .992 .980 .958 .76$\pm$.04
-TopKAvg500 .992 .993 .995 .995 .993 .77$\pm$.04
-TopKAvg1000 .992 .994 .996 .998 .997 .75$\pm$.03
-CombTopK5_200 1.00 .998 .998 .983 .962 1.32$\pm$.06
-CombTopK50_500 1.00 1.00 1.00 .999 .996 1.85$\pm$.03
-Books BruteForce 1.00 1.00 1.00 1.00 1.00 149.02$\pm$.49
-TopKPerEmbd5 .907 .915 .809 .509 .396 20.17$\pm$.06
-TopKPerEmbd10 .936 .986 .914 .698 .561 20.33$\pm$.06
-TopKPerEmbd50 .993 .992 .994 .956 .902 21.09$\pm$.04
-TopKPerEmbd100 1.00 .995 .996 .985 .959 22.53$\pm$.06
-TopKAvg200 1.00 .978 .948 .845 .767 .68$\pm$.03
-TopKAvg500 1.00 .992 .996 .919 .875 .68$\pm$.03
-TopKAvg1000 1.00 1.00 .996 .963 .939 .84$\pm$.01
-TopKAvg2000 1.00 1.00 1.00 .980 .968 1.04$\pm$.02
-TopKAvg4000 1.00 1.00 1.00 .996 .987 1.64$\pm$.04
-CombTopK5_200 .979 .984 .981 .892 .818 20.29$\pm$.05
-CombTopK50_500 .993 .989 .994 .975 .961 21.57$\pm$.06
-CombTopK100_1000 1.00 .997 .996 .993 .992 23.00$\pm$.08
-Table 4: Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency with standard deviation measured over the evaluation of retrieval for a batch of queries (where the batch size is 32).
+ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 2.73$\pm$0.01
+TopKPerEmbd5 0.69 0.67 0.63 0.46 0.40 1.61$\pm$0.05
+TopKPerEmbd10 0.95 0.88 0.82 0.65 0.57 1.65$\pm$0.06
+TopKPerEmbd50 1.00 1.00 0.99 0.95 0.92 1.64$\pm$0.05
+TopKPerEmbd100 1.00 1.00 1.00 0.99 0.98 2.31$\pm$0.02
+TopKAvg200 1.00 1.00 1.00 0.99 0.97 1.19$\pm$0.04
+TopKAvg500 1.00 1.00 1.00 1.00 1.00 1.22$\pm$0.05
+CombTopK5_200 1.00 1.00 1.00 1.00 1.00 1.86$\pm$0.06
+Books BruteForce 1.00 1.00 1.00 1.00 1.00 128.36$\pm$0.30
+TopKPerEmbd5 1.00 0.92 0.90 0.61 0.48 20.47$\pm$0.07
+TopKPerEmbd50 1.00 1.00 1.00 0.98 0.93 21.60$\pm$0.08
+TopKPerEmbd100 1.00 1.00 1.00 0.99 0.97 23.32$\pm$0.08
+TopKPerEmbd200 1.00 1.00 1.00 1.00 1.00 26.55$\pm$0.12
+TopKAvg200 0.97 0.92 0.91 0.76 0.67 1.13$\pm$0.04
+TopKAvg500 0.97 0.98 0.97 0.86 0.81 1.17$\pm$0.04
+TopKAvg1000 0.99 0.98 1.00 0.92 0.88 1.12$\pm$0.05
+TopKAvg2000 1.00 0.99 1.00 0.95 0.92 1.20$\pm$0.02
+TopKAvg4000 1.00 1.00 1.00 0.96 0.95 2.05$\pm$0.01
+TopKAvg8000 1.00 1.00 1.00 0.97 0.97 3.79$\pm$0.01
+CombTopK5_200 1.00 1.00 1.00 0.96 0.95 20.75$\pm$0.07
+CombTopK50_500 1.00 1.00 1.00 0.99 0.96 22.12$\pm$0.07
+CombTopK100_1000 1.00 1.00 1.00 0.99 0.98 24.02$\pm$0.13
+CombTopK200_2000 1.00 1.00 1.00 1.00 1.00 28.01$\pm$0.11
+NQ320K BruteForce 1.00 1.00 1.00 1.00 1.00 37.74$\pm$.47
+TopKPerEmbd5 1.00 1.00 1.00 0.96 1.00 4.71$\pm$0.08
+TopKPerEmbd10 1.00 1.00 1.00 0.98 1.00 4.83$\pm$0.08
+TopKPerEmbd50 1.00 1.00 1.00 1.00 1.00 6.31$\pm$0.09
+TopKAvg100 1.00 1.00 1.00 1.00 1.00 0.57$\pm$0.05
+CombTopK5_100 1.00 1.00 1.00 1.00 1.00 5.28$\pm$0.08
+Table 5.
+Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency measured over a batch of queries (where the batch size is 32).
 (Relative) hit rate higher than .99 is marked in bold.
-We evaluate the following methods on their top $K$ retrieval performance:
+We evaluate the following methods for top $K$ retrieval performance:
   • Brute-force top $K$ (BruteForce): Evaluate the MoL scores for all items and return the top $K$ items.
@@
 […]
+For each dataset, we evaluate top $K$ retrieval methods based on the best performing model configurations reported in Table 3 and Table 4.
 Table 5 shows the hit rate (HR) and latency of all the methods.
 The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$.
-We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms [19, 4, 42].
-We set batch size to $32$ for ML-1M, ML-20M, and Books.
-We perform evaluation on a single NVIDIA RTX 6000 Ada GPU.
+We measure latency by evaluating a batch of 32 retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023).
+We omit ML-1M as its size is small (Table 2).
+We set batch size to $32$ for all datasets.
+We perform evaluation on a single RTX 6000 Ada GPU.
 We report latency averaged over 20 warm runs.
-We observe that our approximate heuristics achieve high HR with oversampling.
-For example, TopKAvg500 is >.$>.99$ in relative HR across the board for ML-1M and ML-20M, and TopKAvg4000 is >.$>.99$ in relative HR across the board for Books.
+We observe that our approximate heuristics achieve high HR with overfetching.
+For example, TopKAvg500 achieves >.$>.99$ in relative HR across the board for ML-20M, and TopKAvg100 achieves >.$>.99$ in relative HR across the board for NQ320K.
 In addition, the combined top $K$ algorithm can outperform both TopKPerEmbd and TopKAvg of the corresponding configurations, sometimes significantly, e.g., CombTopK5_200 vs.
 TopKPerEmbd5 and TopKAvg200 on Books.
-This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions vary in MoL.
-In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baseline, especially as the number of items in the dataset becomes large.
-For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of 4×4$\times$4 × and 91×91$\times$91 × in latency for ML-20M and Books, respectively.
-While the latency of the retrieval grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods.
-For example, the latency of retrieval increases by 4.8×4.8$\times$4.8 × from ML-1M to ML-20M in BruteForce, while the growth rate is 1.8×1.8$\times$1.8 × and $\times1$ for TopKPerEmbd100 and TopKAvg1000 respectively.
-Thus, we expect that the speedup of the approximate methods to become even more prominent as the size of the dataset further increases.
-We also notice that TopKAvg tends to be more efficient than TopKPerEmbd with comparable HR, e.g., TopKAvg2000 vs.
-TopKPerEmbd100 on Books with $\times2$ speedup in latency.
-We believe that this is mainly due to two reasons.
-First, when the HR is comparable, the maximal number of candidate items from TopKPerEmbd is larger than that of TopKAvg.
-Second, compared to TopKPerEmbd, the computation of TopKAvg is agnostic to the number of embedding sets because of the materialization optimization described in Section 3.2.
-Interestingly, we also see that the combined top $K$ is more efficient than the summation of the latency of its individual components, e.g., CombTopK5_200 is 1.5×1.5$\times$1.5 × faster than the sum of the latency from TopKPerEmbd5 and TopKAvg200 on ML-20M.
-This is because our implementation reduces the overhead of the combined method by consolidating some shared processing of the two components.
-Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ can be used if the target HR is extremely stringent.
-5 Related work
+This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions, $\pi_{p}(q,x)$s, vary in MoL.
+In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baselines, especially as the number of items in the dataset becomes large.
+For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of $\times6$ in latency for NQ320K.
+While the algorithm latency grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods.
+For example, the algorithm latency increases by 47.0×47.0$\times$47.0 × from ML-20M to Books in BruteForce, while the growth rate is 10.1×10.1$\times$10.1 × and 1.0×1.0$\times$1.0 × for TopKPerEmbd100 and TopKAvg500, respectively.
+Thus, we expect the speedup of the approximate methods to become even more pronounced with larger datasets.
+We also notice that TopKAvg tends to be more efficient than TopKPerEmbd at comparable HRs, e.g., TopKAvg4000 is 10.5×10.5$\times$10.5 × faster than TopKPerEmbd50 on Books in terms of latency.
+First, the computation in TopKAvg is agnostic to the number of component-level embeddings, $P$, because of the preprocessing described in Section 3.2.
+Second, the branching and deduplication steps in TopKPerEmbd cannot leverage the parallel processing capabilities of GPUs effectively.
+Additionally, the latency of the combined top $K$ algorithm is lower than the sum of its individual components’ latencies; e.g., on ML-20M, CombTopK5_200 has a latency 1.5×1.5$\times$1.5 × lower than the sum of the individual latencies of TopKPerEmbd5 and TopKAvg200.
+This is done by consolidating processing shared by the two components.
+Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ algorithm can be used if the target HR is extremely stringent.
+5. Related work
 Similarity Functions in Retrieval.
-Most information retrieval models in recommendations and natural language processing follow a classical two-stage paradigm [5], where up to billions of items [6, 26, 3] are first filtered down to hundreds in the retrieval stage, followed by the ranking stage that produces the final top 1-10 candidates.
-Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function [5, 21].
+Most information retrieval models in recommendation systems and natural language processing (e.g., question answering) follow a classical two-stage paradigm (Covington et al., 2016; Karpukhin et al., 2020), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019; Zhai et al., 2023; Borisyuk et al., 2024) are first filtered down to hundreds in the retrieval stage, followed by another stage (e.g., ranking in recommendation systems or generation in RAG (Lewis et al., 2020)) that produces the final results.
+Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function (Covington et al., 2016; Karpukhin et al., 2020; Ni et al., 2022a, b).
 Researchers quickly realized that dot products limited retrieval stage’s performance, and explored various learned similarity-based approaches.
-Prominent variants include maximum similarity based on multiple embeddings [26, 22, 34], specialized neural networks, often based on Hadamard products [14, 37, 39, 3], and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal [16, 45, 46] or explicitly in the “generative retrieval” setups [8, 38, 40, 36].
-It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendations [33] and DSI in NLP [36].
-Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items [3].
-Despite these challenges, significant gains (17% gains at Hit Rate@100 [42] to 24% gains at Hit Rate@400 [3]) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions [34, 42], implicit diversification done as part of beam search [8, 36], explicit representation of side-information using special neural architectures [3], and hardware-aware similarity function and inference algorithm design on GPUs [4, 42, 3].
-Efficient Nearest Neighbor Search (NNS).
-Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases.
+Prominent variants include maximum similarity based on multiple embeddings (Li et al., 2019; Khattab and Zaharia, 2020; Santhanam et al., 2022), specialized neural networks, often leveraging Hadamard products (He et al., 2017; Tan et al., 2020; Wang et al., 2022a; Borisyuk et al., 2024), and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal (Jasinska et al., 2016; Zhu et al., 2018; Zhuo et al., 2020) or explicitly in the “generative retrieval” setups (De Cao et al., 2021; Tay et al., 2022; Wang et al., 2022b; Bevilacqua et al., 2022; Zhuang et al., 2023; Sun et al., 2023).
+It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendation systems (Rendle et al., 2020) and DSI for QA scenarios in NLP (Sun et al., 2023).
+Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items (Eksombatchai et al., 2018; Zhai et al., 2023; Borisyuk et al., 2024).
+Despite these challenges, significant gains (17% gains at Hit Rate@100 (Zhai et al., 2023) to 24% gains at Hit Rate@400 (Borisyuk et al., 2024)) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions (Santhanam et al., 2022; Zhai et al., 2023), implicit diversification done as part of beam search (Gao et al., 2021), explicit incorporation of side-information using special neural architectures (Zhai et al., 2023; Borisyuk et al., 2024), and hardware-aware similarity function and inference algorithm design on GPUs (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024; Borisyuk et al., 2024).
+Load Balancing for Conditional Computations in Neural Networks
+Conditional computations have been widely utilized in deep learning models (Bengio et al., 2016; Shazeer et al., 2017; Chen et al., 2023).
+Regularization losses have been proposed based on the observation that an ideal policy should evenly utilize all compute units in aggregate while being sparse at an individual example level (Bengio et al., 2016).
+Mixture-of-experts, a common way to implement conditional computations, has been widely used in language and vision domains (Chen et al., 2023; Shazeer et al., 2017) where mutual information-based regularization losses between experts and tasks (Chen et al., 2023) and experts and tokens (Shen et al., 2023) have been shown to help with various architectures.
+Efficient Nearest Neighbor Search (NNS)
+NNS has been a popular research topic due to their critical role in large-scale retrieval and vector databases.
 Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS).
-Various techniques were proposed and analyzed, including tree structures [1, 32], locality sensitive hashing [10, 35], production quantization [18, 11], data partitioning [27, 44], graph-based methods [28, 17], and so on.
-The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results [46].
-Algorithms based on implicit [28, 17, 39, 31] or explicit graphs [39] have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case [30].
-Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques [35, 11] no longer fully utilize GPUs; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work [4, 42, 31, 3].
-6 Conclusion
-We have analyzed techniques for efficient retrieval with learned similarities in this work.
-We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and further empirically learnable – MoL consistently outperforms dot products in HR across 6 scenarios by an average of 24.2% in HR@1 and 14.3% in HR@10, setting new state-of-the-art on recommendation datasets.
-We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds.
-Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads.
+Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011; Guo et al., 2016), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on.
+The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020).
+Algorithms based on implicit (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019; Wang et al., 2022a; Ootomo et al., 2024) or explicit graphs (Wang et al., 2022a) have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case (Morozov and Babenko, 2018).
+Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize the compute; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed recently (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024; Borisyuk et al., 2024).
+6. Conclusion
+We have analyzed techniques for efficient retrieval with expressive learned similarities in this work.
+We begin by showing Mixture-of-Logits (MoL) is a universal approximator of all similarity functions, and further empirically learnable – MoL with our proposed load balancing loss consistently outperforms dot products (dense retrieval), sparse retrieval, and generative retrieval approaches across Recommendation Systems and Question Answering scenarios, setting new state-of-the-art across common, heterogeneous benchmark datasets.
+We next propose exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and error bounds.
+Our approximate top $K$ algorithms can reach >.$>.99$ of Hit Rate relative to exact algorithms, while achieving up to $\times6$ reduction in end-to-end latency and with minimal indexing overheads.
 We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations.
-Given MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items [42, 3], our work concludes both theoretical and practical justifications for migrating vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.
-Limitations
-[…]
+Given MoL’s impressive empirical performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items (Zhai et al., 2023; Borisyuk et al., 2024) and broad applicability across heterogeneous scenarios, our work provides strong theoretical and practical justifications for migrating web-scale vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.
 Appendix A Experiment Setups
-Reproducibility.
-Our code is publicly available online, and detailed implementations and hyperparameter settings for reproducing our experiment results can be found at the following url: https://github.com/bailuding/rails.
-We discuss key settings below.
-Parameterization of low-rank (“component-level”) embeddings, $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$.
-We note that embedding parameterization was not explicitly discussed in the original Mixture-of-Logits paper [42].
-A specific implementation utilizing cluster id information has been proposed and discussed in a follow up work [3], where in the context of large-scale recommendation systems, the authors suggested cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$.
-However, we cannot easily access similar information in the publicly available MovieLens and Amazon Reviews datasets.
-We therefore follow implementation provided by the authors [42] and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems [24].
-All other $f_{p}(q)$s and $g_{p}(x)$s are obtained by applying an MLP on top of query-side/item-side representations in standard sequential recommendation setups [15, 20].
-Parameterization of $\pi_{p}(q,x)$ matrices.
-We follow the implementation provided by the authors [42], which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU [7] non-linearity.
-The inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the learned low-rank embeddings.
-Hyperparameter settings.
+A.1. Reproducibility
+The implementations and hyperparameter settings for reproducing our experiment results can be found at https://github.com/bailuding/rails. We discuss specific details below.
+A.2. Parameterization of low-rank (“component-level”) embeddings
+In this section, we elaborate on the embedding parameterization methods for MoL that we discussed in Section 2.2.
+A.2.1. Recommendation Systems
+Prior work have shown that careful parameterization of low-rank (“component-level”) embeddings, or $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$, can significantly improve MoL’s performance (Borisyuk et al., 2024).
+In the context of large-scale recommendation systems, cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$ (Borisyuk et al., 2024).
+However, we cannot easily access similar information in the publicly available MovieLens (Harper and Konstan, 2015) and Amazon Reviews (McAuley et al., 2015) datasets.
+We therefore follow implementation provided by (Zhai et al., 2023) and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems (Koren et al., 2009) that we discussed in Section 2.2.
+All other component-level embeddings, $f_{p}(q)$s and $g_{p}(x)$s, are obtained by applying a multi-layer perceptron (MLP) on top of query-side/item-side representations in standard sequential recommendation setups (Hidasi et al., 2016; Kang and McAuley, 2018).
+The overall setup is illustrated on the right hand side of Figure 3.
+Figure 3.
+Illustration of how to parameterize the embeddings to adapt Mixture-of-logits (MoL) learned similarity to various retrieval scenarios, with a language model (LM) finetuning use case in question answering (characterized by a single homogeneous feature) shown on the left, and a recommendation systems use case (characterized by a large number of heterogeneous features) shown on the right.
+For the Question Answering example on the left, $SP_{1},\ldots,SP_{N}$ represents the original SentencePiece (Kudo and Richardson, 2018) tokens that are inputs to the pre-trained language model LM, e.g., T5 (Raffel et al., 2023).
+$Q_{1},Q_{2},\ldots,Q_{P_{Q}}$ and $X_{1},X_{2},\ldots,X_{P_{X}}$ represent the special aggregation tokens we add to the LM tokenizer for pooling information across the sequence.
+The “Parameterized Pooling” component uses a $D$-dimensional embedding as input to parameterize, at an example-level, how to weight each of the (max_seq_len) encoder outputs for the $P_{Q}$/$P_{X}$ MoL component-level embeddings.
+A.2.2. Question Answering (QA)
+Unlike Recommendation Systems, retrieval models used in question answering generally take the full semantic representation(s) of the query and/or the document as input, and are finetuned on top of pre-trained language models with homogeneous inputs, or wordpiece / sentencepiece tokens.
+Our MoL embedding construction consists of two components, special aggregation tokens and parameterized pooling.
+We present embedding construction on the query side first.
+Special Aggregation Tokens
+Given both queries and documents are represented as token sequences (e.g., SentencePieces (Kudo and Richardson, 2018) in T5 (Raffel et al., 2023)), we propose to add special tokens that can be used to aggregate different aspects of information as part of the overall self-attention based language model.
+Specifically, on the query side, let the tokenized sequence be $SP_{1},SP_{2},\ldots,SP_{N}$.
+During finetuning of the pretrained language model, we create $P_{Q}$ special tokens, $Q_{1},\ldots,Q_{P_{Q}}$, and add them to the vocabulary of the query tokenizer.
+We also append those exact same $P_{Q}$ tokens before $SP_{1},SP_{2},\ldots,SP_{N}$, so that the $P_{Q}$ special tokens can be used to aggregate information across the query input using early-fusion mechanisms.
+Note that many question answering scenarios (Karpukhin et al., 2020; De Cao et al., 2021; Wang et al., 2022b; Ni et al., 2022b; Sun et al., 2023) utilize bidirectional language models for retrieval, like BERT (Devlin et al., 2019) or T5 (Raffel et al., 2023); for recent unidirectional language models, we can add the special aggregation tokens $X_{1},\ldots,X_{P_{X}}$ and $Q_{1},\ldots,Q_{P_{Q}}$ to the end of the input sequence instead.
+Our construction can also be viewed as a way to extend the CLS token in BERT (Devlin et al., 2019; Chang et al., 2023) to cover multiple aspects of information, in a way that encourages diversity via the $\mathcal{L}_{MI}$ load balancing loss discussed in Section 2.
+Parameterized Pooling
+We next add a pooling layer after the language model to encourage learning of aggregation mechanisms separate from language semantics.
+For each position $1\leq p\leq P_{Q}$, this pooling layer defines a probability distribution over different positions in language model’s outputs, or $(0,\ldots,max\_seq\_len-1)$.
+We further parameterize the pooling layer, using the $D$-dimensional embedding at the first position after encoders.
+This enables us to define a pooling policy, at an example-level, how to weight each of the $max\_seq\_len$ LM encoder outputs to arrive at the $P_{Q}$ MoL embeddings.
+The embedding construction on the item-side is identical. We illustrate the overall finetuning setup we use for question answering on the left hand side of Figure 3.
+A.3. Parameterization of $\pi_{p}(q,x)$ matrices
+We follow the implementation provided in the original MoL paper (Zhai et al., 2023), which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU (Elfwing et al., 2017) non-linearity.
+For recommendation datasets (ML-1M, ML-20M, Books), the inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the low-rank embeddings.
+For question answering datasets (NQ320K), we only use the last part – the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the low-rank embeddings – as inputs to this MLP.
+A.4. Hyperparameter settings
+A.4.1. Recommendation Systems
 We use an identical number of sampled negatives for dot product baselines (cosine similarity, “SASRec”, “HSTU” rows in Table 3) and Mixture-of-Logits (“SASRec + MoL”, “HSTU + MoL” rows in Table 3) to ensure a fair comparison, which is $128$ for ML-1M and ML-20M and $512$ for Amazon Books following prior work.
-For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$), whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets.
-Due to resource constraints, we followed initial hyperparameters provided by the authors [42] for all other parameters.
-A.1 Efficiency
-Method HR@10 HR@50 MRR Time/Epoch
-ML-1M SASRec .2818 .5470 .1352 37s
-SASRec + MoL .3034 (+7.7%) .5655 (+3.4%) .1448 (+7.1%) 44s (+18.9%)
-HSTU .3332 .5956 .1579 41s
-HSTU + MoL .3460 (+3.8%) .6048 (+1.5%) .1673 (+6.0%) 46s (+12.2%)
-ML-20M SASRec .2883 .5484 .1375 197s
-SASRec + MoL .3124 (+8.4%) .5685 (+3.7%) .1551 (+12.8%) 449s (+127.9%)
-HSTU .3557 .6146 .1800 526s
-HSTU + MoL .3697 (+3.9%) .6244 (+1.6%) .1880 (+4.4%) 773s (+47.0%)
-Books SASRec .0306 .0754 .0153 506s
-SASRec + MoL .0402 (+31.4%) .0970 (+28.6%) .0203 (+32.7%) 964s (+90.5%)
-HSTU .0469 .1066 .0233 1883s
-HSTU + MoL .0613 (+30.7%) .1292 (+21.2%) .0315 (+35.2%) 2236s (+18.7%)
-Table 5: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task.
-Note that these results are identical to those reported in Table 3 but with relative gains (MoL vs dot product) and training cost changes illustrated under different sequential encoder backbones.
-Training efficiency.
-We measure the training efficiency of the setups in Section 4 over one epoch of data, utilizing a single RTX 6000 Ada GPU and with a constant batch size of 128 for ML-1M and ML-20M, and 64 for Amazon Books in Table 5.
-Consistent with prior findings [42, 3], MoL’s model quality increases do not generally come with a large increase in training time, thanks to MoL’s higher arithmetic intensity on GPUs, with an average increase of 53% across the six scenarios.
-To contextualize this 1.53x training cost, prior work on sequential recommenders, such as self-supervised bidirectional models, came with a 5.9x training time increase on ML-20M and 4.5x training time increase on Amazon Reviews (BERT4Rec vs SASRec rows) [23].
-We expect additional GPU kernel-level optimizations, such as kernel fusions, to further improve efficiency of the “+ MoL” rows.
-Inference efficiency.
-Given the number of items (3K, 24K, and 674K) in these datasets, as discussed in Table 2, these public benchmark datasets are not good candidates for measuring inference latency.
-We note that prior work have reported batched inference latency between 5ms and 20ms on datasets with hundreds of millions to billions of items [42, 3], which is inline with optimized MIPS baselines.
+For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$, whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets.
+We followed initial hyperparameters provided by the authors (Zhai et al., 2023) for all other parameters.
+The models are trained using PyTorch over 1 NVIDIA RTX 6000 Ada GPU for the smaller ML-1M and ML-20M datasets and 2 NVIDIA RTX 6000 Ada GPUs for the larger Books datasets.
+A.4.2. Question Answering (QA)
+We train the model with AdamW optimizer (Loshchilov and Hutter, 2019), and grid searched learning rate in {2e-4, 5e-4, 8e-4} due to the introduction of the parameterized pooling component (Appendix A.2).
+We apply linear scheduling with warm-up over a fixed 10% of the training epochs.
+We train the model on 4 NVIDIA H100 80GB GPUs with a local batch size of 512.
+Note that due to the computational requirements of this dataset, prior work are frequently trained on 8 GPUs (Karpukhin et al., 2020; Wang et al., 2022b) or more, e.g., 32 GPUs in GENRE (De Cao et al., 2021) and 256 TPUs in DSI (Tay et al., 2022).
+We perform in-batch negative sampling, consistent with baselines (Karpukhin et al., 2020; Ni et al., 2022a).
+For MoL hyperparameters, we grid searched $P_{Q}$ and $P_{X}$ in {(2, 2), (4, 4), (8, 8), (16, 16)}, kept $d_{P}$ identical to the embedding dimension of the pretrained language model ($768$), and selected the best hyperparameters utilizing a validation set.
+Appendix B Examples for exact and approximate top $K$ retrieval algorithms
+We provide examples for our exact- and approximate- retrieval algorithms discussed in Section 3 to facilitate understanding.
+Retrieval with Exact Algorithm (Section 3.1).
+item $\langle f_{1},g_{1}\rangle$ $\langle f_{2},g_{2}\rangle$ $\pi_{1}$ $\pi_{2}$ $\phi$
+a 1 1 0.5 0.5 1.0
+b 0.8 0 0.5 0.5 0.4
+c 0 0.8 0.5 0.5 0.4
+d 0.7 0 1 0 0.7
+e 0.2 0.2 0.5 0.5 0.2
+Table 6.
+Example illustrating how exact- and approximate- top-$k$ algorithms work.
+We consider a fixed query $q$, and provide inner products $\langle f_{p}(x),g_{p}(q)\rangle$s, gating weights $\pi_{p}$s, and learned similarity scores $\phi$s for that query.
+Table 6 shows an example with $4$ items and $2$ pairs (groups) of embeddings ($P=2$).
+Assume the goal is to retrieve the top $K=2$ items.
+In the first stage, we retrieve the top $2$ items for each embedding set, i.e., item $a,b,c$ are retrieved.
+For each of the retrieved items, we calculate the MoL scores based on their gating weights, i.e., $1.0,0.4,0.4$ for $a,b$, and $c$, respectively.
+Here, $S_{min}$ is set to $0.4$.
+In the second stage, we retrieve all items with $\langle f,g\rangle\geq 0.4$ for each embedding set, i.e., item $d$, and then calculate their corresponding MoL score, i.e., $0.7$.
+The algorithm returns $a$ and $d$ as the top $2$ items.
+Retrieval with Approximate Algorithm (Theorem 3).
+Consider again the example shown in Table 6.
+Assume we want to retrieve the top $2$ items.
+With top $K$ over component-level embeddings, item $a,b,c$ are retrieved, and $S_{k}=0.4$.
+Since $S$ is calculated as $\max\{0.7,0.2\}$, the upper bound of the gap is $0.3$.
+Here, the gap is exact, i.e., the actual second largest MoL score with item $d$ is $0.3$ higher than the second largest MoL score from the set of retrieved items $\{a,b,c\}$.
+If we bound the gap with the retrieved items only, i.e., $a,b,c$, then we will get a looser bound of $0.8-0.4=0.4$.